### v2

# DBudgetKV: Dynamic Budget in KV Cache Compression for Ensuring Optimal Performance

Abstract
To alleviate memory burden during inference of large language models (LLMs), numerous studies have focused on compressing the KV cache by exploring aspects such as attention sparsity. These techniques are often designed with a pre-defined KV budget; however, as the optimal budget varies by different input lengths and task types, the existence of a fixed budget could result in inconsistent performance accepting inputs of diverse domains. To address this limitation, we propose a new KV cache compression objective: to always ensure the full-cache performance regardless of specific inputs, while maximizing KV cache pruning as much as possible. To achieve this goal, we introduce a novel KV cache compression method dubbed DBudgetKV, which features an attention-based metric to signal when the remaining KV cache is unlikely to match the full-cache performance, then halting the pruning process. Empirical evaluation spanning diverse context lengths, task types, and model sizes suggests that our method achieves lossless KV pruning effectively and robustly, exceeding 25% compression ratio on average. Furthermore, our method is easy to integrate within LLM inference, not only optimizing memory space, but also showing reduced inference time compared to existing methods.
1 Introduction
Most large language models (LLMs) follows autoregressive generation, where they rely on KV Cache to store intermediate states during inference [1, 2, 3]. Specifically, LLMs access the cached key and value vectors of past tokens during the attention calculation, thereby speeding up inference by re-using these KV cache [4, 5, 6]. However, when the model size and the input length increase, the memory required to maintain the KV cache also grows proportionally. Since these vectors are usually stored within GPU memory, managing the KV cache efficiently has become crucial to mitigate the overall memory consumption and computational overhead during LLM inference. For instance, Llama3 8B model [7] requires 1 GB of KV cache memory for 2K-token inputs, while its 70B counterpart demands a gigantic memory up to 50GB for 20K tokens.
Consequently, recent works have proposed to effectively reduce KV cache, leveraging the sparsity of the attention mechanism: certain positions are more pivotal during the generation process, while those less important ones could be pruned with minimal performance degradation; thus, the full KV cache is not always necessary. These observations have led to several optimization methods, such as H2O [8], ScissorHands [9], StreamingLLM [10], and FastGen [11], where they aim to identify and retain only the most salient token positions, discarding less critical ones based on certain pruning criteria. Apart from pruning, other paradigms have been proposed as well to reduce KV consumption, such as KVMerger [12] and D2O [13] that merge or compress KV vectors instead of hard-pruning.
While these studies have successfully demonstrated that a decent portion of KV cache can be discarded in practice, there remains one common issue that has yet to be addressed: these pruning techniques typically require a pre-defined KV cache budget, of which its optimal threshold could vary significantly according to specific tasks or inputs. In real-world scenarios, one would need to carefully tune budget thresholds optimized for different domains; otherwise, setting a uniform threshold in diverse domains may potentially bring large degradation on certain tasks. We reckon the need for a pre-defined threshold as an overlooked limitation that can lead to unstable and inconsistent performance in practice.
Table 1: With pre-defined cache budgets, KV pruning methods on Llama3 exhibit inconsistent performance across domains (percentage relative to full-cache scores).
Methods GSM8K GPQA CoQA
Budget = 90%
H2O 99.1% 96.9% 100.2%
SLM 96.7% 99.2% 100.2%
SnapKV 93.3% 96.1% 99.9%
Budget = 20%
H2O 3.2% 8.5% 96.5%
SLM 4.3% 11.5% 76.1%
SnapKV 1.7% 3.9% 98.6%
To further illustrate, Table 1 shows that when retaining 90% KV cache, Llama3-8B-Instruct could achieve near full-cache performance across three domains. However, the performance is shown inconsistent with 20% budget; especially on the math benchmark GSM8K [14], all three KV pruning methods drastically fall to less than 5% relative to the full-cache performance.
To address the aforementioned limitation, in this work, we propose a new objective for KV cache compression: our method aims to strike towards full-cache performance regardless of specific inputs, while maximizing KV cache pruning as much as possible. Towards this objective, we propose an adaptive pruning method, dubbed DBudgetKV, which Dynamically adjusts KV cache retention per input, without the need of pre-defined memory Budget. In particular, DBudgetKV automatically identifies input-dependent KV cache pruning that leads to minimal performance degradation. A key implication is that the method tends to realize a higher compression ratio for simpler tasks, while allocating more cache resources for complicated tasks.
Our method features a two-step process: (1) after the prefilling stage, rank all input tokens according to an importance metric; (2) discard the KV cache of each token sequentially, until hitting a stopping criterion. In choosing the importance metric and the stopping criterion, DBudgetKV takes into account two important factors from previous works: i) position bias, where tokens at the beginning and the end are generally more significant than those in the middle [10, 15]; ii) attention score, where tokens receiving higher scores usually contribute more information to subsequent generation [1, 2, 8].
To eliminate the pre-defined KV budget while always targeting full-cache performance, the stopping criterion becomes crucial, such that it should signal whenever the remaining KV cache is unlikely to match the full-cache capacity, thereby fulfilling the objective to approximate full performance with dynamic cache pruning. In DBudgetKV, we design a simple attention-based metric as the stopping condition: the metric compares the difference between the Frobenius form of the full attention matrix and the reduced attention matrix of which pruned positions’ scores are set to zero. When the difference exceeds a universal threshold, the pruning process halts. In practice, we find that this metric empirically correlates well with the full-cache performance regardless of input types, serving as an estimator to bridge the pruning state and the subsequent generation.
Figure 1: The overall workflow of DBudgetKV in Section 3.2. After prefilling, tokens are initially ranked based on their positions, followed by the eviction of the least significant tokens (per layer), whose halting condition is determined by the norm value of the reduced attention matrix. The KV cache for the remaining tokens are then preserved to subsequent generation.
To examine the performance consistency across diverse domains, our experiments adopt 13 datasets varying diverse context lengths and tasks, e.g. mathematical and commonsense reasoning, reading comprehension and coding, demonstrating that DBudgetKV achieves our objective effectively and robustly. With dynamic KV budgets, the resulting inference is on par or even surpasses the full-cache performance with multiple LLMs of different sizes, including Llama3 [7], Qwen2.5 [16], and Mistral [17]. The dynamic pruning of DBudgetKV automatically allocates relatively high cache budget on math benchmarks to prevent degradation, while achieving high compression ratio on comprehension tasks by up to 85%. With Llama3-8B, DBudgetKV reaches an average budget ratio of 63.7% with minimal degradation across inputs. Overall, empirical results suggest that our proposed method is able to accomplish lossless KV cache pruning with strong generalizability. Besides, it does not rely on specific model architectures, and is orthogonal to other KV optimization techniques. We further conduct experiments for in-depth analysis and ablation studies, where we validate that each step of DBudgetKV is necessary in achieving the desired performance, depicting the rationality of various design choices and hyper-parameters. Our contributions can be summarized as follows:
  • To the best of our knowledge, we are the first to call for the objective that ensures full-cache performance while maximizing the KV cache pruning, addressing the limitation of inconsistent performance with fixed KV budget thresholds.
  • We introduce DBudgetKV that employs a straightforward two-step process, where the dynamic pruning is halted with minimal degradation to the generation performance.
  • Extensive evaluation with diverse tasks and models indicates that compared to prior works, our method achieves lossless KV cache compression effectively, efficiently and robustly.
2 Related Work
KV Cache Compression
Recent advancements in KV cache compression have emerged to address memory constraints for model inference. Scissorhands [9] observes the repetitive nature of attention patterns and the persistence of token importance. It then preserves the most important tokens based on attention scores to achieve KV cache compression. H2O [8] introduces dynamic eviction policies that strategically balance retention of recent tokens and historically significant "heavy hitters". StreamingLLM [10] enables infinite-length sequence processing without fine-tuning by stabilizing attention through landmark tokens. SnapKV [18] enhances compression efficiency through attention score-based selection and clustering of critical KV positions. FastGen [11] proposes an adaptive KV cache management system that customizes retention strategies per attention head. Other paradigms beyond KV pruning have been proposed along the way, such as KV cache merging [19].
As prior KV pruning methods typically require setting a pre-defined memory budget, they are by design inherently unable to ensure consistent performance across various domains and datasets, which could prevent their widespread deployment in practice. Our method targets to lift such constraint, achieving near full-cache performance and dynamic budget size without the need of additional tuning.
Attention Pattern
The study of attention head specialization has paved the way for optimizing memory and computation in Transformer models. Early research on encoder architectures shows that some attention heads have persistent positional biases, like consistently focusing on adjacent tokens, indicating redundancy in their learned patterns. Rho-1 [20] reveals that the importance of individual heads varies significantly across tasks, suggesting that not all attention computations are equally crucial during inference, and reveal the potential for dynamically pruning or approximating redundant computations. Recent works [6, 21] have shifted the focus to decoder-only models, where the quadratic memory overhead of retaining intermediate KV vectors poses a significant challenge. Focused Transformer [22] extends the context length of LLMs through a contrastive learning-inspired training process. Landmark-Attention [23] uses a “landmark” token to represent each block of the input and trains the attention to use it for selecting relevant blocks, enabling retrieval of blocks directly through the attention mechanism instead of by relying on a separate mechanism.
Our work leverages attention’s structural regularities, particularly the predictable, position-specific behaviors of certain heads, to dynamically reduce the memory footprint of cached KV states. By identifying and exploiting the properties with position bias and attention distribution, we enable efficient inference without compromising model performance, aligning with the broader goal of dynamic KV cache pruning for various inputs.
3 Methodology
We first elaborate our objective in Sec. 3.1 as a new direction for KV cache compression. Distinct from previous works, we aim to compress the KV cache as much as possible while ensuring LLMs’ performance remains intact. We then delineate our proposed approach along with its motivation behind in Sec. 3.2. Important implementation details are next discussed in Sec. 3.3.
3.1 Objective of KV Cache Compression
As mentioned in Sec. 1, most previous works on KV cache pruning require a pre-defined memory budget beforehand, either by fixing the number of cached positions directly [24, 18], or retaining positions by a fixed ratio of input length [8, 13]. While these approaches are effective in certain circumstances, they may encounter practical challenges when deployed in real-world scenarios.
It is forthcoming that the optimal budgets for LLM inference in real-world scenarios are evidently infeasible to enumerate, as they could vary across tasks and domains, especially for open-domain instructions. For example, mathematical tasks are typically concise and require inference based on all given conditions, thus many positions may logically contribute to the generation, necessitating a relatively higher memory budget. Conversely, for article reading comprehension, LLM may only need a small set of critical positions for the answer generation, featuring a smaller budget.
As a side effect, inference with fixed KV cache pruning may experience performance instability across different inputs, which can be observed by our analysis in Sec. 4.2. We regard it as a practical limitation that motivates the inception of this work.
In addressing this limitation, we resort to a dynamic paradigm that eliminates the need for manually setting a memory budget. We introduce a new objective for KV cache compression, aiming to automatically reduce the KV cache as much as possible, while fulfilling the condition that the method should always strike towards full-cache performance. In this way, one could utilize such pruning method off-the-shelf, without spending time and resources on tuning usage-specific hyper-parameters, which we hope would further advance the research development of KV cache compression techniques.
3.2 DBudgetKV
Towards our objective, we design a novel ranking-based pruning method applied after the prefilling stage, comprising two steps. First, all tokens are ranked according to an importance metric per Transformers layer. Then, the system keeps removing the KV cache sequentially, until hitting a stopping criterion. Ultimately, our proposed approach does not need to fix an optimal budget for storing KV cache; rather, it relies on the stopping criterion to determine dynamically, which should work uniformly regardless of the inputs.
The motivation behind our two-step procedure comes from two phenomena brought by previous works. First, there exists position bias such that positions in the beginning and the end are typically important for the subsequent generation, as observed by StreamingLLM [10]. Second, attention scores are usually an effective indicator of the token importance, as leveraged by many pruning techniques [9, 8, 18]. For our objective, we propose to combine the two important properties in our approach, described in details as follows.
Importance Ranking
Denote a LLM input sequence as $X=\{x_{1},x_{2},\dots,x_{n}\}$, where each Transformers layer originally consists of $n$ positions of KV vectors per attention head. As the pruning process starts from the least important KV vectors, all tokens are firstly ranked according to an importance strategy.
To this end, we resort to a simple ranking strategy leveraging the position bias, without undergoing any additional computation. As identified by StreamingLLM, a significant portion of attentions is allocated to the initial tokens, regardless of their relevance to the language modeling task, known as attention sinks. Inspired by this, we employ a purely position-based importance evaluation strategy: we assume that the first $m$ tokens are always crucial; while among the remaining $n-m$ tokens, those towards the end are more significant. By this strategy, the importance in $X$ is ranked from the highest to the lowest as $\{x_{1},x_{2},\dots,x_{m},x_{n},x_{n-1},\dots,x_{m+1}\}$, denoted as $\widehat{X}$, where $m$ is a chosen hyperparameter that works regardless of specific input sequences.
Like previous KV compression works, our choice on importance ranking is empirically supported rather than theoretically based. We have also experimented other attention-based importance ranking, which are provided in Sec. 4.5. The choice of position-based ranking is empirically shown both empirically effective and superior on computational overhead.
Stopping Criterion
With the initially ranked KV vectors, our method then sequentially decides whether to continue or terminate pruning, which requires an evaluation metric as the stopping criterion. Since one evidently cannot know the precise impact of a KV cache position before the generation, we adopt a metric that empirically correlates well with the performance change when discarding a position, serving as a bridge to ensure a minimal degradation upon the full cache performance.
Inspired by Devoto et al. [25], we utilize the Frobenius norm of the attention matrix that fits the requirement well empirically. For a matrix $M\in\mathbb{R}^{m\times n}$, the Frobenius norm of $M$ is the matrix-version of the L2 norm, expressed as $||M||_{F}=\sqrt{\sum_{i=1}^{m}\sum_{j=1}^{n}|M_{i,j}|^{2}}$.
Figure 2: Preliminary experiments on the relationship between the Frobenius norm difference (in percentage) obtained from randomly masking tokens, and the corresponding model performance on GSM8K.
We conduct preliminary experiments to validate its reliability. We employ Llama2-7B and randomly mask a certain portion of the KV cache, setting the corresponding values in the attention matrix $A\in\mathbb{R}^{n\times n}$ to zero. As illustrated in Figure 2, there is a certain degree of positive correlation between its Frobenius norm and the model’s performance, which are shown generalized across different domains. The simple Frobenius norm works empirically well in subsequent large-scale evaluation in Sec. 4.
As the input sequence length $n$ increases, the time and space overhead for norm calculation also grows significantly by $O(n^{2})$. Hence, we next aim to reduce the computational scale from the naive norm calculation. We observe that the Frobenius norm of the last row of $A$ also correlates with the model’s performance; equivalently speaking, the attention distribution of the latest token is good enough to serve as an estimator. Formally, we utilize the last $k$ rows (latest $k$ tokens) of $A$ and reduce to a single attention vector $A^{\prime}\in\mathbb{R}^{1\times n}$. The score $s_{j}$ for a position $j\in[1,n]$ in $A^{\prime}$ is:
$s_{j}=\frac{\sum_{i=k}^{n}A_{i,j}}{\sum_{i=k}^{n}\mathbf{1}_{\{A_{i,j}\neq 0\}}}$ (1)
We then use the Frobenius norm of $A^{\prime}$, equivalently its L2 norm, as the evaluation metric, which reduces the overall complexity to $O(n)$.
Finally, for each position in the ranking order, DBudgetKV compares the difference between the Frobenius norm of the original attention matrix $A$, and the Frobenius norm of the reduced attention vector $A^{\prime}$ of which the pruned positions are set to 00. We set a universal threshold as 1%, such that when the new norm exceeds 1% difference from the original norm, it hits the stopping criterion. The KV vectors of the remaining positions are then all kept, while the preceding positions have their KV cache discarded. In this way, our method achieves dynamic KV cache pruning while targeting to maintain the full-cache performance, assisted by the norm metric.
3.3 Implemetation Details
Though our proposed pruning process is conceptually sequential, it is implemented efficiently by PyTorch’s operators, such that the stopping positions of all layers are identified directly in parallel without any looping operations. Empirical analysis in Sec. 4.4 shows that our proposed method achieves higher efficiency compared to three popular baselines.
Specifically, we reorder $A^{\prime}$ based on Token Importance Ranking $\widehat{X}$ to obtain $A^{\prime}_{sort}$, where $A^{\prime}_{sort}[0]$ corresponds to the least important token and $A^{\prime}_{sort}[-1]$ corresponds to the most important token. Next, we compute the cumulative square-sum of each element in $A^{\prime}_{sort}$:
$A^{\prime}_{cumsum}[i]=\sqrt{\sum_{k=i}^{n}|A^{\prime}_{sort}[k]|^{2}}$ (2)
such that $A^{\prime}_{cumsum}[i]$ represents the Frobenius Norm of the attention matrix after removing all tokens to the left of $i^{th}$ token. We then divide each element in $A^{\prime}_{cumsum}$ by $||A^{\prime}||_{F}$ and subtract this value from 1 to determine the difference between each new matrix and $A^{\prime}$. The torch where operation allows us to directly identify the position of the token that satisfies the condition of not exceeding a 1% difference, ultimately yielding the set of tokens for which the KV cache is called to retain. The final kv cache compression algorithm of DBudgetKV is presented in Algorithm 1. Since PyTorch allows specifying the dimension for operations, DBudgetKV can run in parallel across multiple attention heads.
Retaining Bottom Layers
Additionally, we have discovered that applying DBudgetKV to the model’s 0th and 1st layers results in poor performance when the model operates under a high budget. We hypothesize that this may be due to the relatively uniform distribution of attention scores in these initial layers, leading to the premature discarding of tokens that are crucial for subsequent layers, ultimately causing the entire model’s inference to collapse. A more detailed analysis of this issue is provided in Appendix C. Based on this finding, we initiate the KV cache compression operations starting from the model’s 2nd layer. Regarding ablation studies are further conducted in Sec.4.5.
4 Experiments
4.1 Experimental Settings
Table 2: Performance of DBudgetKV and its comparison with three KV cache pruning models on 13 datasets. Bold numbers indicate the best results aside from full-cache. Italics represent the real budget utilized by DBudgetKV. The correspondence between abbreviations and their full names of tasks and datasets can be found in Appendix. Avg. calculates the mean ratio of the model’s performance using different KV cache compression methods to its performance with the full cache. All average results (except for budget) are adjusted by subtracting 1 to provide a more intuitive understanding of the effectiveness of different pruning methods.
  Methods Math & Science CR Single-Doc QA Multi-Doc QA Summarization FSL Code   Avg.
GSM8K
GPQA
TheoQA
ThQA
CoQA
NrtvQA
Qasper
2WkMQA
Musique
QMSum
M-News
TriviaQA
Lcc
Llama3-8B-Instruct Full 75.28 29.02 21.29 25.59 52.74 24.06 43.91 35.33 14.77 22.27 27.37 70.26 19.16 100%
Oursk=1 76.50 30.13 23.03 26.09 52.86 23.44 37.38 36.21 15.96 21.95 27.88 64.24 19.31 +0.12%
Budget 93.2% 86.7% 92.8% 78.7% 81.5% 48.7% 46.4% 45.8% 43.7% 15.0% 76.4% 41.0% 78.0% 63.68%
H2O0.9 74.60 28.13 21.69 25.41 52.83 24.55 41.73 33.95 15.23 22.51 27.64 69.77 19.13 -0.39%
SLM0.9 72.78 28.79 20.75 25.15 52.83 23.75 43.63 32.68 15.86 22.48 27.21 69.97 19.66 -0.59%
SnapKV0.9 70.20 27.90 20.08 25.38 52.72 24.04 43.93 33.92 15.56 22.49 27.65 70.18 19.05 -1.12%
H2O0.5 31.08 2.90 16.47 17.99 42.37 23.14 43.05 32.79 15.77 22.79 26.24 69.83 19.18 -17.63%
SLM0.5 3.41 6.03 8.84 18.75 44.68 20.94 37.73 31.20 15.29 21.80 25.93 67.79 19.12 -24.73%
SnapKV0.5 12.96 7.81 16.87 19.52 42.84 23.40 43.93 33.98 15.94 22.67 26.07 69.66 19.19 -17.03%
H2O0.2 2.43 2.46 6.56 20.67 50.91 23.54 42.00 32.50 15.67 22.16 23.68 69.88 19.13 -23.33%
SLM0.2 3.21 3.35 8.82 18.02 40.11 20.12 36.45 29.91 15.01 21.11 24.68 66.17 18.69 -28.21%
SnapKV0.2 1.29 1.12 5.89 20.36 52.00 23.10 42.78 32.98 17.01 22.45 24.14 69.63 19.34 -22.89%
Mistral-7B-Instruct Full 33.36 29.24 6.83 20.82 39.68 28.74 37.80 33.87 22.88 22.19 22.94 86.87 16.08 100%
Oursk=1 31.54 29.02 6.71 20.81 39.80 27.20 38.93 33.46 22.15 22.39 22.67 86.81 15.33 -1.50%
Budget 89.6% 90.5% 84.2% 92.3% 84.1% 78.0% 97.9% 86.7% 74.4% 84.3% 89.1% 87.2% 89.4% 86.75%
H2O0.9 19.18 24.11 6.96 20.61 39.74 27.36 38.38 35.23 23.24 22.13 23.41 86.46 16.01 -4.29%
SLM0.9 31.16 27.68 5.89 19.88 39.15 27.26 37.66 35.06 21.94 22.31 21.73 86.63 13.67 -4.44%
SnapKV0.9 27.60 25.67 7.10 20.23 39.76 27.43 38.27 35.66 22.99 22.16 23.31 86.46 16.07 -1.90%
H2O0.5 2.50 8.26 5.89 20.28 38.53 27.62 36.97 34.31 23.17 22.32 22.60 86.52 16.45 -14.31%
SLM0.5 2.43 6.47 1.14 18.27 32.76 24.91 33.80 32.73 20.68 21.48 18.54 86.45 13.58 -27.61%
SnapKV0.5 3.03 8.93 6.83 19.03 39.22 27.15 38.27 34.60 22.94 21.93 22.68 86.31 16.18 -13.41%
H2O0.2 1.52 4.46 4.02 18.45 35.55 27.28 33.45 34.36 23.21 21.80 20.93 85.99 14.82 -23.55%
SLM0.2 1.21 1.79 0.54 18.35 25.45 24.89 28.39 29.53 16.74 20.60 16.75 80.12 14.27 -35.58%
SnapKV0.2 1.44 0.20 3.35 17.81 37.44 26.76 36.02 33.90 23.05 22.13 20.88 84.27 16.20 -24.28%
Qwen2.5-7B-Instruct Full 88.02 31.70 29.85 24.55 61.43 20.81 43.17 47.15 30.70 23.64 24.24 87.64 2.44 100%
Oursk=1 88.02 31.25 30.39 24.48 60.01 20.66 42.74 47.20 29.56 22.64 24.04 87.65 3.58 +2.63%
Budget 90.7% 88.8% 86.5% 69.2% 84.1% 65.1% 69.2% 73.3% 64.4% 70.6% 85.4% 56.6% 84.4% 76.02%
H2O0.9 83.47 26.56 30.25 24.40 61.29 20.85 43.26 47.90 30.73 23.42 24.35 87.57 2.45 -1.46%
SLM0.9 88.93 30.80 28.25 24.30 60.91 19.86 42.54 46.02 28.75 23.06 23.42 87.03 3.05 -0.41%
SnapKV0.9 80.14 30.13 28.11 24.30 61.26 20.90 43.31 47.81 30.69 23.90 24.29 87.57 2.55 -1.01%
H2O0.5 34.12 14.73 20.88 23.10 59.69 21.59 43.37 47.60 30.81 21.00 22.95 85.54 2.24 -13.47%
SLM0.5 3.26 1.34 10.58 23.75 49.87 19.29 36.45 42.42 25.33 21.24 22.77 87.33 3.40 -23.56%
SnapKV0.5 20.77 12.28 22.76 22.25 60.21 20.93 43.22 47.45 30.61 23.76 22.70 87.57 2.21 -14.39%
H2O0.2 5.91 3.57 17.40 21.70 55.39 21.41 40.30 45.84 29.72 23.04 20.84 87.99 2.15 -13.89%
SLM0.2 1.06 4.24 4.82 23.85 40.92 18.17 29.22 39.27 21.71 20.32 18.94 77.80 1.46 -31.66%
SnapKV0.2 2.58 7.37 15.13 21.52 57.06 21.40 40.93 46.96 29.87 23.58 20.69 87.10 2.44 -13.26%
Backbones
Our experiments are conducted with three LLMs: Llama3-Instruct with size of 8B/70B, Mistral-7B-Instruct-V0.3, and Qwen2.5-Instruct with size of 7B/32B/72B. We implement DBudgetKV upon the codebase of SnapKV [^1] https://github.com/FasterDecoding/SnapKV. For the reduced attention matrix $A^{\prime}$, we set $k=1$ in practice (ablation provided in Sec. 4.5), and set $m=4$ for importance ranking.
Datasets
For comprehensive evaluation, we evaluate DBudgetKV using datasets of both short and long context length. For tasks of relatively short inputs, the models are assessed across mathematics, science, and commonsense reasoning on multiple datasets including GSM8K [14], GPQA [26], TheoremQA [27], TruthfulQA [28], and CoQA [29]. For long-context tasks, we adopt tasks and datasets from Longbench [30]. Appendix A provides a detailed description and statistics of these datasets along with how they are utilized in experiments.
Evaluation Protocol
Our primary objective is to evaluate the capability of DBudgetKV to compress KV cache while striking for a full-cache performance. To this end, we compare DBudgetKV with the full KV cache performance and calculate the average compression ratio per dataset. Additionally, we select three previous KV cache compression methods with fixed budget size, including: Heavy Hitter Oracle (H2O) [8], StreamingLLM (SLM) [10], and SnapKV [18]. We compare with three budgets for these three methods at 90%, 50% and 20% compression ratio respectively.
Apart from the main experiments, we further evaluate various design factors of DBudgetKV via ablation studies, such as different strategies of token ranking importance, the selection of $k$ for the attention matrix, and the necessity to retain bottom layers. Finally, we also provide efficiency analysis on the time and space overhead comparing against other KV cache pruning methods in Sec. 4.4.
4.2 Main Results
The main results are shown in Table 2 according to our evaluation protocol, where DBudgetKV demonstrates strong advantages with robust performance across diverse tasks and models:
Table 3: Performance of LLMs with scales of 70B, 32B, and 72B across five datasets using DBudgetKV.
GSM8K CoQA NQA Musique QMSum
Llama3-70B-Instruct
Full 89.69 60.36 27.15 29.31 22.52
Oursk=1 89.76 60.38 26.94 28.88 22.30
Budget 93.3% 77.2% 58.7% 60.8% 52.9%
Qwen2.5-32B-Instruct
Full 91.36 58.03 24.78 40.04 22.84
Ours 91.81 57.29 22.65 40.54 22.52
Budget 91.6% 85.0% 68.2% 74.4% 76.8%
Qwen2.5-72B-Instruct
Full 90.22 54.14 24.36 42.13 23.93
Ours 90.30 54.19 24.10 41.70 23.31
Budget 90.9% 87.6% 63.4% 65.0% 68.6%
• DBudgetKV is able to fulfill our objective, capable of performing near-lossless dynamic compression across different models, varying input lengths, and diverse task types. Interestingly, with Llama3-8B and Qwen2.5-7B, DBudgetKV even surpasses the full-cache performance by 0.12% and 2.63% respectively, utilizing an average of 63.68% and 76.02% KV cache. With Mistral, DBudgetKV also manages to achieve nearly 20% KV cache compression with only a 1.5% performance reduction. These results indicate that our proposed method can be suited for real-world scenarios with no bother by domain-specific budget thresholds. In stark contrast, previous methods achieve a consistent full-cache performance only when manually determined a high budget ratio, e.g. 90%. However, when the budget is reduced, e.g. 50% and 20%, the degradation can become severe on certain datasets, distinct from DBudgetKV that automatically adjusts the pruning to always prioritize full performance.
• DBudgetKV naturally reflects the difficulty of the generation task. As in Table 2, the dynamic budget ratio is high on Math&Science datasets (over 90%), while much lower on QA or Summarization datasets (as low as 15%). This observation is in line with our intuition, where inference on concise but hard tasks, such as math problems, requires more context and more precise calculation, resulting in higher budget allocation. Thus, our method design serves beyond for inference efficiency, but could also be leveraged for input analysis in a broader scope.
• Besides the dynamic compression, DBudgetKV also outperforms the three 90%-budget baselines, while itself uses less than 90% budget, and achieves the highest score in 20 out of 39 comparisons. This suggests that by taking into account both the position bias and attention scores, our method accompolishes more effective pruning than the baselines that only consider one aspect, highlighting the importance of integrating both dimensions for effective KV cache compression.
It is worth noting that the scope of this work is not to propose yet another KV cache pruning method. Instead, we identify the common limitation of previous KV pruning methods, and call for optimal KV pruning without a pre-defined budget. Consequently, the consistent performance across domains after KV pruning is the ultimate goal of our method, instead of achieving the highest compression ratio on certain datasets. Under our objective, we leave the further improvement on compression ratio as a promising future research direction.
4.3 DBudgetKV’s Generalization Ability
We next conduct experiments to examine whether our design and hyper-parameters can be generalized to LLMs of larger sizes. On five datasets in Table 3, DBudgetKV with Llama3-70B and Qwen2.5-32B/72B demonstrates consistent near full-cache performance with the same DBudgetKV setting. Notably, the averaged compression ratio increases on datasets of longer context, achieving nearly a 50% compression. But most importantly, they are able to maintain lossless performance across inputs, validating the robustness and versatility for its general applicability.
4.4 Efficiency Analysis
Table 4: The average inference time and pruning time of Llama3 with size of 8B/70B across six datasets, measured in seconds, with lower values indicating better performance.
GSM8K CoQA NarrativeQA Musique QMSum TriviaQA Avg.
Overall Prune Overall Prune Overall Prune Overall Prune Overall Prune Overall Prune Overall Prune
Llama3-8B-Instruct
Full 4.693 0.289 3.441 3.659 5.458 2.717 3.376
Oursk=1 4.638 0.034 0.331 0.033 3.159 0.255 3.658 0.264 5.330 0.241 2.684 0.211 3.300 0.173
H2O0.5 4.686 0.020 0.290 0.018 3.243 0.264 3.679 0.266 5.398 0.249 2.612 0.237 3.318 0.176
SLM0.5 6.071 0.006 0.301 0.030 3.230 0.257 3.784 0.259 5.454 0.234 2.628 0.208 3.578 0.166
SnapKV0.5 5.874 0.021 0.285 0.019 3.307 0.265 3.721 0.266 5.519 0.264 2.650 0.265 3.559 0.183
Llama3-70B-Instruct
Full 17.504 1.167 6.285 6.565 13.993 5.302 8.469
Oursk=1 15.975 0.154 1.253 0.290 6.042 2.345 7.412 2.352 13.835 3.028 5.224 1.602 8.290 1.623
H2O0.5 19.062 0.114 1.059 0.240 6.788 2.298 7.082 2.307 14.360 3.712 5.566 1.308 8.986 1.663
SLM0.5 22.079 0.150 1.138 0.229 6.804 2.069 7.107 2.053 14.734 3.700 5.576 1.367 9.573 1.595
SnapKV0.5 16.517 0.117 1.116 0.241 6.795 2.474 7.085 2.468 13.846 2.793 5.562 1.083 8.487 1.529
Table 5: Performance comparison for ablation studies in Section 4.5, with different $k$ values for attention matrix reduction, ranking methods and universal thresholds.
Methods GSM8K CoQA NQA Musique QMSum
Full 75.28 52.74 24.06 14.77 22.27
Post=1% Performance Using Different $k$
$k=1$ 76.50 52.86 23.44 15.96 21.95
Budget 93.2% 81.5% 48.7% 43.7% 15.0%
$k=1\%n$ 76.19 52.87 23.17 15.40 21.94
Budget 95.1% 94.8% 77.3% 77.2% 59.5%
$k=5\%n$ 75.59 52.75 21.62 13.71 21.43
Budget 98.6% 96.2% 45.1% 30.0% 33.1%
$k=10\%n$ 76.72 52.85 9.74 13.67 21.11
Budget 96.3% 96.8% 32.0% 19.7% 29.3%
$k=1$ Performance with Different Ranking Method
Attnt=1% 2.35 43.68 13.37 9.58 17.62
Budget 20.0% 9.7% 6.4% 6.4% 6.4%
Attnt=0.01% 54.66 51.58 17.45 14.79 20.90
Budget 61.7% 28.6% 6.7% 7.9% 7.1%
Posk=1 Performance with Different Threshold $t$
${t=0.1\%}$ 76.35 52.84 22.73 17.04 21.90
Budget 99.2% 97.0% 77.2% 85.0% 73.9%
${t=10\%}$ 26.23 49.69 18.35 15.16 18.81
Budget 55.8% 37.2% 15.27% 18.5% 6.5%
In this section, a quantitative study is conducted on whether DBudgetKV improves inference time, apart from the space reduction from KV cache pruning. We compare the time usage on six datasets with Llama3-8B/70B using DBudgetKV, along with three baselines with the 50% budget setting. Table 4 reports the average generation time after the prefilling stage of each sample (Overall), as well as the time needed to complete pruning (Prune) before the generation.
From the results, DBudgetKV achieves the best performance in 8 out of 12 comparisons of overall generation time, and achieves the best average time for both 8B and 70B LLMs, suggesting that although DBudgetKV requires a slightly longer time for pruning itself compared to three other approaches, the total generation speed of DBudgetKV is evidently advantageous. It also performs consistently across models of different scales, underscoring its time efficiency.
4.5 Ablation Study
The ablation study is conducted to explore the impact of various configurations of DBudgetKV. We select Llama-3-8B-Instruct and perform experiments across five datasets presented in Table 5. Appendix B presents additional results and analysis from the ablation study.
Attention Matrix Reduction
The reduced attention matrix $A^{\prime}$ in Sec. 3.2 aggregates attention scores from the last $k$ positions. The upper part of Table 5 illustrates the model’s performance and the actual budget when setting $k=1$, $1\%n$, $5\%n$, and $10\%n$ ($n$ being the number of input tokens). It is clear that setting $k$ as $1$ achieves a significantly reduced budget, thus a higher compression ratio, with almost no change in performance compared to $1\%n$ and $5\%n$. On the other hand, while $10\%n$ can compress more KV cache, it fails to maintain performance (for instance, on NarrativeQA, the former achieves a performance of 9.74 using 32% of the budget, whereas the latter scores 21.44 using 48.7% of the budget). What’s even better is that since $k=1$ requires the least amount of computation, relying solely on the scores from the last token, the complexity of obtaining $A^{\prime}$ becomes $O(1)$, independent of the sequence length. The advantages of both high efficacy and low overhead make $k=1$ a solid design choice for calculating the norm metric.
Importance Ranking Strategies
We investigate DBudgetKV potentials through ranking by each token’s average attention score received from other tokens, similar to previous approaches such as H2O. The results are shown in the middle of Table 5: under the same experimental settings, the attention-based ranking struggles to maintain the model’s performance. Therefore, the position-based strategy is more suited for our goal.
Table 6: Performance comparison between DBudgetKV and StreamingLLM with the first two layers frozen (SLMF) under difference KV cache budgets ($0.9$, $0.5$ and $0.2$).
GSM8K GPQA CoQA NrtvQA QMSum TriviaQA
Full 75.28 29.02 52.74 24.06 22.27 70.26
Oursk=1 76.50 30.13 52.86 23.44 21.95 64.24
Budget 93.2% 86.7% 81.5% 48.7% 15.0% 41.0%
SLM0.9 72.78 28.79 52.83 23.75 22.48 69.97
SLMF0.9 76.42 29.24 52.89 22.95 22.47 69.10
SLM0.5 3.41 6.03 44.68 20.94 21.80 67.79
SLMF0.5 4.55 3.13 44.59 21.31 21.88 68.04
SLM0.2 3.21 3.35 40.11 20.12 21.11 66.17
SLMF0.2 1.21 1.12 38.44 19.52 20.61 65.69
Attention Norm Threshold
As we adopt the universal threshold as 1% for the attention norm difference, we comprehensively study the effects of smaller or larger thresholds, as shown in the bottom part of Table 5. Intuitively, a larger threshold allows for a higher compression ratio, while a smaller threshold does the opposite. The conclusion is clear that when the threshold is set to a smaller value, such as 0.1%, the average budget increases as expected, yet the model’s performance sees little improvement. Conversely, when the threshold is set to 10%, it prunes more KV cache, but the model’s performance significantly deteriorates. Thus, we deem 1% as a reasonable threshold for universal uses across models and tasks.
Comparison with StreamingLLM
We perform additional experiments on six datasets using StreamingLLM (with Llama3-8B) that has the first two layers frozen, as an ablation study on the effects of the proposed pruning and halting process. As shown in Table 6, despite having the first two layers frozen, StreamingLLM still falls short of achieving the objective proposed in this work, not able to remain optimal consistently across datasets like GSM8K and GPQA as the budget varies. It is demonstrated that solely leveraging the position bias as in StreamingLLM is not able to maintain the full performance; the proposed pruning and halting mechanism that integrates the attention norm metric is necessary in achieving the goal for lossless pruning.
5 Conclusion
In this study, we introduce a novel KV cache compression objective designed to approximate the full-cache performance, independent of specific inputs, while optimizing resource utilization through targeted KV cache pruning. Our approach, termed DBudgetKV, employs a straightforward yet effective two-step process for each layer of LLMs. Tokens are initially ranked solely based on their positions, followed by the eviction of the least significant tokens as determined by the norm value of the reduced attention matrix. Comprehensive experiments conducted across diverse datasets, encompassing a variety of tasks and context lengths, demonstrate that DBudgetKV achieves nearly lossless compression, while with notable space and time reduction, successfully fulfilling our objective for greater practical values.
Appendix A Datasets Used in Experiments
In this section, we provide a comprehensive overview of all the tasks and datasets utilized in the experiments in this paper.
Math & Science
This task evaluates the model’s ability to tackle mathematical and scientific problems. By directly inputting questions and comparing the model’s output with the correct answers, we calculate the model’s Accuracy on these datasets: GSM8K is a dataset for evaluating model’s math-solving skills, featuring 8,000 elementary-level math word problems requiring basic arithmetic and reasoning. GPQA tests model’s understanding of physics concepts and problem-solving across various topics, assessing scientific reasoning abilities. TheoremQA evaluates model’s grasp and application of mathematical theorems, ranging from simple applications to complex proofs, testing advanced math skills.
Commonsense Reasoning (CR)
This task evaluates model’s ability to make deductions and understand everyday situations using implicit knowledge and logical inference. TruthfulQA (ThQA) evaluates model’s ability to generate accurate and truthful responses, testing models on distinguishing fact from fiction, especially in areas prone to misconceptions. We use BLEU as the metric. CoQA assesses model’s ability to understand and respond to questions in a conversational context, focusing on maintaining coherence and context throughout a dialogue. We use F1 Score as the metric.
Single Document QA (Single-Doc QA)
This task assesses the model’s reading comprehension skills when dealing with a single, extended document. NarrativeQA [31] is a dataset designed to evaluate model’s ability to comprehend and answer questions based on narrative texts, focusing on understanding stories and their underlying themes. Qasper [32] is a dataset aimed at assessing model’s capability to extract and answer questions from academic papers, emphasizing understanding complex scientific information. We employ F1 Score as the metric for above two datasets.
Multi-Document QA (Multi-Doc QA)
This task evaluates the model’s reading comprehension capabilities across multiple extended documents. 2WikiMultiHopQA (2WKMQA) [33] is a dataset designed to test model’s ability to perform multi-hop reasoning and answer complex questions using information from multiple Wikipedia articles. MuSiQue [34] evaluates model’s skill in integrating and reasoning over information from multiple sources to answer comprehensive questions accurately. We leverage F1 Score as the metric for above two datasets.
Summarization
This task examines the model’s ability to comprehend and summarize lengthy documents. QMSum [35] is a dataset for evaluating model’s ability to generate concise summaries of meeting transcripts, focusing on capturing the key points from multi-party discussions. Multi-News (M-News) [36] is a dataset that challenges models to create coherent summaries by synthesizing information from multiple news articles on the same topic. We use Rouge-L as the metric for above two datasets.
Few-Shot Learning (FSL)
This task assesses the model’s few-shot learning capabilities. TriviaQA [37] is a dataset designed to assess model’s ability to retrieve and answer questions based on large collections of trivia, emphasizing comprehension and factual recall. We use F1 Score as the metric.
Code
This task evaluates the model’s ability to complete and generate code. LCC [38] is a dataset focused on evaluating models’ ability to understand and generate code by considering extended code contexts, enhancing the ability to reason over complex programming structures. We use Edit Sim as the metric.
Appendix B Ablation
Table 7: Performance comparison with different $k$, expanded from Table 5.
Methods GSM8K CoQA NQA Musique QMSum
Full 75.28 52.74 24.06 14.77 22.27
Post=1% Performance Using Different $k$
$k=1$ 76.50 52.86 23.44 15.96 21.95
Budget 93.2% 81.5% 48.7% 43.7% 15.0%
$k=1\%n$ 76.19 52.87 23.17 15.40 21.94
Budget 95.1% 94.8% 77.3% 77.2% 59.5%
$k=2\%n$ 76.19 52.82 22.87 14.34 21.74
Budget 96.0% 96.1% 69.8% 62.0% 47.1%
$k=3\%n$ 75.82 52.80 23.03 13.61 21.56
Budget 98.6% 95.7% 60.3% 46.0% 39.5%
$k=4\%n$ 75.21 52.76 22.96 13.45 21.47
Budget 98.8% 95.8% 51.9% 35.2% 35.4%
$k=5\%n$ 75.59 52.75 21.62 13.71 21.43
Budget 98.6% 96.2% 45.1% 30.0% 33.1%
$k=6\%n$ 75.66 52.81 21.95 14.33 20.94
Budget 98.5% 96.6% 40.4% 25.8% 31.6%
$k=7\%n$ 76.57 52.88 21.18 13.66 20.94
Budget 98.0% 96.7% 37.3% 22.5% 30.6%
$k=8\%n$ 76.35 52.86 20.36 13.75 21.11
Budget 97.4% 96.7% 35.1% 20.9% 30.1%
$k=9\%n$ 76.65 52.84 19.99 13.66 20.76
Budget 96.8% 96.8% 33.4% 19.9% 29.5%
$k=10\%n$ 76.72 52.85 9.74 13.67 21.11
Budget 96.3% 96.8% 32.0% 19.7% 29.3%
In this section, we present additional ablation study results for Section 4.5. By setting various values for $k$, we expand upon the results shown in Table 5. The expanded results are displayed in Table 7. These experiments facilitate a deeper understanding of how different parameter settings impact model performance and provide a basis for optimizing parameter selection.
As shown in Table 7, setting $k=1$ not only conserves pruning time but also achieves better model performance with a reduced budget.
Appendix C Freeze the first two layers of LLM
In this section, we demonstrate the necessity of using DBudgetKV to compress the KV cache starting from the 2nd layer of LLMs. We apply DBudgetKV on Llama2-7B-Chat and conduct experiments on GSM8K and CoQA. We first perform a case study, followed by a comparison of the effects of not freezing layers of the model on its performance.
As shown in Table 9, when the threshold is set to 1% and no layers are frozen, the outputs for two examples on GSM8K and CoQA are incorrect and lack logical coherence. However, the budget exceeds 90% in both cases. By observing the actual budget of each layer, we can see that for these two examples, the budgets for layer 0 and 1 are relatively low, while the budgets from layer 2 to layer 30 are the same and very high, with the last layer being low again. We analyze that since the model, in the first two layers, is not yet able to identify truly important tokens, the attention distribution is relatively uniform. This may lead to the model discarding truly important tokens during eviction, resulting in subsequent generations being unable to access this information, ultimately causing the output to fail.
Table 8: Performance on GSM8K and CoQA using DBudgetKV with different frozen layers.
Datasets Layer None 0 0,1 0,1,2 0,1,2,3 0,1,31
GSM8K Oursk=1 0.014 0.250 0.264 0.252 0.264 0.252
Budget 93.1% 97.0% 95.4% 97.5% 97.7% 98.2%
CoQA Oursk=1 1.47 52.80 53.58 53.58 53.32 53.46
Budget 92.7% 94.7% 95.5% 98.3% 98.4% 99.2%
Based on the above case study, we attempt to freeze certain layers of the model and explore the optimal freezing configuration. We continue to test the results of freezing different layers. The resuls are shown in Table 8. From the table, we can observe that freezing the first two layers achieves a balance between model performance and budget. Moreover, freezing the 31st layer does not significantly enhance the model’s performance and instead leads to an increase in budget. Consequently, DBudgetKV ultimately opts to begin KV cache compression from the 2nd layer of the model.
Appendix D Full Algorithm of DBudgetKV
  Input: Prompt, Threshold $t$
  Output: Compressed KV Cache
  Create Empty List $K_{c},V_{c}$
  for Transformer Layer $L_{i}$ in LLM do
     $Q^{i},K^{i},V^{i}\leftarrow L_{i}(\text{Prompt})$
     $R^{i}\leftarrow$ Postion-Based Importance Rank
     $A_{last}^{i}\leftarrow\operatorname{Attention}(Q^{i}[\dots,-1,:],K^{iT})$
     $F_{b}^{i}\leftarrow\operatorname{Frobenius}(A_{last}^{i})$
     $A_{last}^{i}\leftarrow\operatorname{Square}(A_{last}^{i})$
     Reorder $A_{last}^{i}$ by Rank $R$
     $A_{cumsum}^{i}\leftarrow\operatorname{Cumsum}(A_{last}^{i})$
     $A_{cumsum}^{i}\leftarrow\operatorname{Sqrt}(A_{cumsum}^{i})$
     $A_{ratio}^{i}\leftarrow(F_{b}^{i}-A_{cumsum}^{i})/F_{b}^{i}$
     $\text{Index}^{i}\leftarrow\operatorname{Max}(\operatorname{Where}(A_{ratio}^{i}<=t))$
     $K_{c}^{i}\leftarrow\operatorname{Compress}K^{i}\text{by}R^{i}[I:]$
     $V_{c}^{i}\leftarrow\operatorname{Compress}V^{i}\text{by}R^{i}[I:]$
     Append $K_{c}^{i},V_{c}^{i}$ to $K_{c},V_{c}$
  end for
  return  $K_{c},V_{c}$
Algorithm 1 DBudgetKV
Table 9: Case Study.
GSM8K
Input
A robe takes 2 bolts of blue fiber a-
nd half that much white fiber. How
many bolts in total does it take?
Budget
Layer 0: 67.71
Layer 1: 82.29
Layer 2$\sim$30: 95.83
Layer 31: 37.50
Avg.: 93.90
Output
I have determined by answering the
answer to the format of bolts bolts b-
olts…(repeat)
Ground-Truth 3
CoQA
Input
You are given a story and a question.
Answer the question as concisely as
you can…Question: What color was
Cotton?
Budget
Layer 0: 42.14
Layer 1: 42.54
Layer 2$\sim$30: 99.19
Layer 31: 79.64
Avg.: 95.03
Output
Question: Question: What is the que-
stion: What is the question:
Groud-Truth White
Appendix E Limitations
The main limitation of DBudgetKV stems from the gap between its current compression budget and the true optimal budget. This can be seen in Table 2 that for certain scenarios, e.g. QMSum with Mistral-7B, DBudgetKV reaches 84.3% budget while 50% budget is also viable with no performance degradation. We regard this gap as room for improvement in future work, facilitating more aggressive KV cache compression while ensuring the full-cache performance.
Another limitation of DBudgetKV is that though it demonstrates almost lossless compression, we rely on empirical experiments and there is no hard guarantee on the full-performance constraint. In Table 2, while both Llama3-8B and Qwen2.5-7B even surpass the full-cache performance, DBudgetKV with Mistral-7B shows trivial degradation of 1.5%. More robust methods could be developed as future work to further strengthen this constraint.
Appendix F Broader Impacts
Our work has made contributions on both a new objective in this line of research, as well as a new method that successfully fulfills this objective. Since all current KV cache pruning methods require a pre-defined KV budget threshold, they either have inconsistent and unstable performance across different input lengths, domains and tasks, or they need careful tuning of their budgets per dataset to ensure the full-cache performance. Our proposed method evaluated on 13 diverse datasets, including short and long-context QA, coding and math, demonstrates its effectiveness on addressing this limitation, which automatically adjusts the budget to reach full-cache performance regardless of the inputs.