### v3

# Towards Threshold-Free KV Cache Pruning

Abstract
To reduce memory consumption during LLM inference, prior works have proposed numerous methods that focus on KV cache pruning based on various criteria. While these techniques often accomplish lossless memory reduction on many datasets, they often rely on an under-emphasized condition: a dataset/domain-specific budget size threshold needs to be pre-determined to achieve the optimal performance. However, such input-specific tuning may be considerably limited in real-world scenarios, as open-domain inputs span diverse domains, lengths and difficulty levels, without clear boundaries for pre-tuning. Thus, the dependence of an input-sensitive threshold can be an inherent limitation that may cause large degradation on arbitrary inputs.
In this work, we propose a new objective that lifts the threshold constraints for robust KV pruning, calling for “threshold-free” methods that automatically adjust budget sizes while ensuring full-cache performance. We then propose a novel method ReFreeKV as the first solution fulfilling this objective, validated by intensive experiments on 13 datasets of diverse context lengths, task types, and model sizes.
Towards Threshold-Free KV Cache Pruning
Xuanfan Nithanks:  Equal contribution. Partial work done during Xuanfan’s internship at WeChat AI.   Liyan Xu11footnotemark: 1333 Project lead: Liyan Xu <liyanlxu@tencent.com>    Chenyang Lyu   Longyue Wang Mo Yu   Lemao Liu   Fandong Meng   Jie Zhou   Piji Li Nanjing University of Aeronautics and Astronautics WeChat AI, Tencent     Independent Researcher xuanfanni@gmail.com   liyanlxu@tencent.com
1 Introduction
As current Transformer-based large language models (LLMs) employs autoregressive generation, they produce KV Cache to store intermediate states during inference DBLP:journals/corr/abs-2307-03109; DBLP:journals/corr/abs-2303-08774; DBLP:journals/corr/abs-2402-06196, where they access the cached key and value vectors of past tokens that typically reside within GPU memory for attention calculation DBLP:conf/nips/VaswaniSPUJGKP17; DBLP:journals/corr/abs-1911-02150; DBLP:conf/emnlp/AinslieLJZLS23. Consequently, managing the KV cache efficiently has become crucial to mitigate the overall memory consumption and computational overhead during LLM inference, as they can become massive and grow proportionally when the model size and the input length increase. For instance, Llama3-8B model DBLP:journals/corr/abs-2407-21783 requires 1GB of KV cache memory for 2K-token inputs, while its 70B counterpart demands a gigantic memory up to 50GB for 20K tokens.
Focusing on this research topic, numerous methods have been recently proposed to effectively reduce the KV footprint after LLM prefilling. Leveraging the sparsity of attention, retaining full KV cache is not always necessary. Several optimization methods, such as H2O DBLP:conf/nips/Zhang00CZC0TRBW23, ScissorHands DBLP:conf/nips/LiuDLWXXKS23, SnapKV DBLP:conf/nips/LiHYVLYCLC24, FastGen DBLP:conf/iclr/Ge0LZ0024, CAKE DBLP:conf/iclr/QinCLHFCLL25, etc., discard the less critical cache positions based on certain pruning criteria. Other paradigms such as KVMerger DBLP:journals/corr/abs-2407-08454 and D2O DBLP:journals/corr/abs-2406-13035 resort to merge or compress KV vectors instead of hard-pruning for achieving the memory reduction effects.
While these studies have confirmed on a number of datasets that a decent portion of KV cache can be pruned, they hinge on a fundamental yet often under-emphasized condition: a threshold, often being dataset-specific, is typically involved for selective tuning to achieve satisfactory results. For instance, D2O takes a pre-defined KV cache budget ratio of 20% to reach full cache performance on LongBench DBLP:conf/acl/BaiLZL0HDLZHDTL24, whereas the required budget jumps to 80% on GSM8K DBLP:journals/corr/abs-2110-14168; likewise, a pre-selected 1024 cache positions is enough for CAKE on LongBench, yet the same budget underperforms on the needle-in-the-haystack test needle.
The existence of such threshold serves fine in idealized settings given dedicated datasets. However, its applicability may be considerably limited in real-world scenarios, where the inputs are intermixed across different domains, lengths and difficulty levels without explicit separation. As a result, optimal thresholds cannot be pre-determined for diverse real-world inputs, potentially making the system less robust and prone to significant performance degradation. For demonstration, Table 1 presents preliminary experiments using H2O, SnapKV and CAKE where inputs from different datasets are mixed to preclude dataset-specific threshold tuning. By varying the KV cache budget ratios, the performance across datasets is shown inconsistent, with a trivial gap under 90% budget but drastically escalating under 20% budget.
Methods GSM8K GPQA CoQA AVG.
Budget = 90%
H2O 99.1% 96.9% 100.2% 98.7%
SnapKV 93.3% 96.1% 99.9% 96.4%
CAKE 98.6% 93.8% 97.0% 96.5%
Budget = 20%
H2O 3.2% 8.5% 96.5% 36.1%
SnapKV 1.7% 3.9% 98.6% 34.7%
CAKE 2.2% 15.4% 96.1% 37.9%
Table 1: KV pruning methods that depend on a preset KV budget threshold can exhibit inconsistent performance across domains (percentage relative to full-cache scores using Llama3), making input-specific threshold selection unavoidable for achieving optimal inference.
In this work, we propose a new objective that lifts the threshold dependency in KV cache pruning, thereby enabling the system to robustly handle arbitrary inputs without input-specific tuning. Specifically, our objective prioritizes two principles: 1) the method operates with a universal threshold insensitive to inputs, effectively rendering it “threshold-free”; and 2) the method should consistently achieve performance comparable to its full-cache counterpart, by automatically adjusting KV cache budgets. Only after satisfying these two criteria should the method pursue the best possible compression ratio. As shown by our full experimental results in Table 2, while prior methods may achieve strong compression on certain datasets, none satisfies both critical aspects by our new objective.
Towards our goal, we then propose a novel method, ReFreeKV, featuring a thReshold-Free KV cache pruning. Our method adopts a two-step process with a novel metric to dynamically control the KV cache budget. Concretely, it first ranks all KV cache positions based on their positional importance, a signal proven useful by StreamingLLM DBLP:conf/iclr/XiaoTCHL24; then, it progressively retains key-value vectors in order, and discards the remaining KV cache once the stopping criterion of our proposed input-insensitive Uni-Metric. For minimal overhead, ReFreeKV is implemented with Torch parallel operators, and our analysis in Section 3.3 shows that its latency is on par with prior efficient KV pruning methods varying batch sizes.
The “threshold-free” aspect stems from the empirical validation of our Uni-Metric described in Section 2.2, whose design is shown to be consistent and insensitive to variations in input domains and sequence lengths. Specifically, it adopts a Universal threshold leveraging the Frobenius norm of attention matrix, such that halting by the same threshold empirically incurs minimal degradation relative to the full-cache capacity. By design, ReFreeKV naturally yields higher compression ratios for simpler tasks, while allocating more cache resources to more complex ones.
Figure 1: The overall workflow of ReFreeKV in Section $2.$ After prefilling, tokens are initially ranked based on their positions, followed by the eviction of the least significant tokens (per layer), whose halting condition is determined by the norm value of the reduced attention matrix. The KV cache for the remaining tokens are then preserved to subsequent generation.
To examine its efficacy, our experiments adopt 13 datasets varying diverse context lengths and tasks, e.g. mathematical, commonsense reasoning, reading comprehension and coding, demonstrating that ReFreeKV fulfills our objective: without an input-specific threshold, the resulting inference is highly comparable or even surpasses the full-cache performance, evaluated with multiple LLMs of different sizes. Using Llama3-8B, it automatically allocates an average KV budget ratio of 63.7%, while retaining 100% full-cache performance across 13 datasets. Extensive experiments and ablation studies empirically demonstrate that ReFreeKV effectively accomplishes our proposed objective. Moreover, as ReFreeKV does not rely on specific model architectures, it has the potentials to further combine with other orthogonal memory optimization paradigms such as quantization DBLP:conf/nips/HooperKMMSKG24; DBLP:journals/corr/abs-2503-18773; DBLP:journals/corr/abs-2502-04420.
2 Methodology
In this section, we first elaborate our motivation and the new objective for KV cache compression that differs from prior works. We then delineate our proposed approach ReFreeKV, along with its key implementation details.
2.1 Threshold-Free: A New Objective
As aforementioned, previous KV cache compression methods require a pre-selected budget threshold. While these methods can perform well on numerous datasets, it is forthcoming that the dependence on a threshold can become an inherent limitation. As the optimal threshold can vary across different inputs, it may not be feasible for such systems to pick appropriate thresholds or maintain stable performance in practical scenarios involving arbitrary inputs and open-domain instructions. Certain inputs may necessitate a relatively higher memory budget, such as tasks with multi-step or mathematical reasoning, whereas certain inputs such as straightforward QA queries need only a small set of KV cache. Real-world inputs, however, are mixed and unpredictable, with no clear boundary by difficulty or domain.
Our propose objective is exactly to remove input-specific threshold constraints, motivating threshold-free methods that ensure consistent performance comparable to full-cache regardless of inputs, obviating the need of tuning for optimal thresholds. Pursuit of the best compression ratio is prioritized only after satisfying the objective. To the best of our knowledge, we are the first to propose an effective solution that fulfills such objective.
2.2 ReFreeKV
ReFreeKV consists of two stages implemented with efficient parallel operators. Conceptually, it first ranks all KV cache positions per layer and per attention head. It then sequentially retains key–value vectors until a stopping condition, determined by our input-insensitive metric with a universal threshold, is met; the KV cache at the remaining positions is subsequently discarded. In line with most prior KV compression works kv_survey, our method is applied once after input prefilling, with the primary goal of reducing memory consumption and an additional benefit of improved throughput (Appendix F).
Initial Ranking
The first stage ranks all KV cache, such that the beginning of the sequence may likely contain more critical information than its latter parts, which forms the basis for downstream sequential eviction.
Building on the positional bias reported in StreamingLLM DBLP:conf/iclr/XiaoTCHL24, we also observe that positions at the beginning and end of the input generally play a more critical role in subsequent generation. We exploit this property in the first stage, and rank KV cache by token positions as follows. Denote a LLM input sequence as $X=\{x_{1},x_{2},\dots,x_{n}\}$, where each Transformers layer originally consists of $n$ positions of KV vectors per attention head. The initial ranking takes the first $m$ positions and reversely takes the remaining nmn-m positions, denoted by $\widehat{X}=\{x_{1},x_{2},\dots,x_{m},x_{n},x_{n-1},\dots,x_{m+1}\}$. Note that $m$ is a chosen hyperparameter that works well regardless of specific input sequences.
Our early experiments also conduct exploration on other ranking strategies, reported in Section 3.4. By empirical validation, the position-based ranking is shown not only effective but also particularly advantageous in terms of computational overhead, constituting the first stage for ReFreeKV. However, pruning solely by positions does not fulfill our objective, as our experiments in Appendix G show that StreamingLLM can exhibit inconsistent performance varying budget sizes, highlighting the need for a more robust eviction strategy.
Eviction by Uni-Metric
With the initial ranking on KV cache, ReFreeKV then sequentially retains KV vectors, and halts upon the stopping condition by an input-insensitive Uni-Metric, after which the remaining cache is effectively evicted.
The design of Uni-Metric is then at the core of this process, which requires to signal the degradation level after removing KV cache of certain positions. we propose a metric that empirically correlates well with the performance change when discarding a position, which could serve as a bridge to ensure a minimal degradation upon the full cache performance. Inspired by DBLP:journals/corr/abs-2406-11430, we utilize the Frobenius norm (L2 norm) of the attention matrix $A\in\mathbb{R}^{n\times n}$ as the Uni-Metric, denoted as $A||_{F}=\sqrt{\sum_{i=1}^{n}\sum_{j=1}^{n}|A_{i,j}|^{2}}$. For each position $i$ in the initially ranked sequence, we compare the Frobenius norm of the original attention matrix, $A||_{F}$, with that of a curated attention matrix, $\widetilde{A_{i}}||_{F}$, in which scores to all positions >i>i in the ranked sequence are masked out, replicating the effect of discarding all KV cache beyond position $i$. Once the norm difference reaches a threshold $T$ at position $i_{\text{prune}}$, the entire process terminates, retaining only the KV cache up to $i_{\text{prune}}$ and discarding the remainder, denoted as:
[…]
The Universal Threshold
To fulfill our objective, the threshold $T$ should ensure near lossless pruning invariant to inputs. Upon empirical search, we identify $T=1\%$ could serve well for this purpose. Figure 2 illustrates how performance varies with changes in the norm across different domains. Preliminary studies indicate that when $T<1\%$, performance remains comparable to the full-cache version robustly. We select $T=1\%$ to balance minimal degradation with maximal cache eviction. The efficacy of Uni-Metric and its universal threshold is validated at full scale in the main experiments presented in Section 3.2 and Figure 3.
Reducing Overhead
As the input sequence length $n$ increases, the time and space overhead for norm calculation on the attention matrix grows by $O(n^{2})$. To reduce the computational scale, we seek to use an approximate norm calculation by $O(n)$. Instead of using the full attention $A$, we reduce $A$ by taking the average of its last $k$ rows to a single attention vector $A^{\prime}\in\mathbb{R}^{1\times n}$. The score $s_{i}$ for a position $i\in[1,n]$ in $A^{\prime}$ is denoted as:
$s_{i}=\frac{\sum_{j=k}^{n}A_{i,j}}{\sum_{j=k}^{n}\mathbf{1}_{\{A_{i,j}\neq 0\}}}$ (2)
Logistics
The rationale behind ReFreeKV ’s two-stage process is that finding the optimal pruning positions in a sequence is a combinatorial problem. By combining ranking with a linear search, the problem becomes tractable. The entire procedure is further optimized through parallel operations, as described as follows.
Figure 2: Performance trends of Llama3-8B, Mistral-7B, and Qwen2.5-7B across varying Uni-Metric thresholds. The x-axis represents the threshold percentage (0.1% to 5%), and the y-axis denotes the performance score normalized by the full-cache performance.
2.3 Implementation Details
As ReFreeKV is conceptually a sequential search process, the design of ReFreeKV allows efficient implementation by PyTorch’s operators, such that the stopping positions of all layers are directly identified in parallel without explicit looping operations. Latency and throughput analyses in Section 3.3 and Appendix F demonstrate that ReFreeKV matches prior popular KV pruning approaches, with negligible latency overhead and improved throughput compared to the baselines.
Specifically, the pruning position $i_{\text{prune}}$ can be determined directly by the combination of Torch cumulative-sum and where operators. We compute the cumulative square-sum of each element in $A^{\prime}_{\text{rank}}$, obtained after the initial ranking on $A^{\prime}$:
[…]
such that $A^{\prime}_{cumsum}[i]$ represents the Frobenius Norm of the attention matrix after removing all cache to the right of position $i$. We then divide $\widetilde{A^{\prime}_{i}}||_{F}$ by $A^{\prime}||_{F}$ to determine the norm difference as in Eq (1). The torch where operation allows us to directly identify the leftmost position that satisfies the $1\%$ universal threshold, ultimately yielding the set of positions for which the KV cache is called to retain. The overall pruning process of ReFreeKV is further presented in Algorithm 1.
Retaining Bottom Layers
In our experiments, we discovered that the LLM’s first two Transformers have a relatively uniform attention distribution, usually requiring to retain most of the cache positions. For simplicity and robustness, we always retain all KV cache of the first two layers in our implementation. More studies on bottom layers are further provided in Appendix B.
Running at Scale
Supporting batch sizes >1>1, ReFreeKV is able to perform the entire pruning process for each sample in parallel. To achieve this, we pad the shorter cache segments and update the attention masks accordingly, allowing the LLM to ignore the padded KV positions. The padding operation has a negligible impact on overall performance. Meanwhile, in popular LLM inference engines, e.g. vLLM kwon2023efficient, it is possible to allocate separate KV cache size for each sample, which aligns well with ReFreeKV. Integration with LLM inference engines such as vLLM is under planning for future development.
3 Experiments
  Methods Math & Science CR Single-Doc QA Multi-Doc QA Summarization FSL Code   Avg.
GSM8K
GPQA
TheoQA
ThQA
CoQA
NrtvQA
Qasper
2WkMQA
Musique
QMSum
M-News
TriviaQA
Lcc
Llama3-8B-Instruct Full 75.28 29.02 21.29 25.59 52.74 24.06 43.91 35.33 14.77 22.27 27.37 70.26 19.16 100%
Oursk=1 76.50 30.13 23.03 26.09 52.86 23.44 37.38 36.21 15.96 21.95 27.88 64.24 19.31 +0.12%
Budget 93.2% 86.7% 92.8% 78.7% 81.5% 48.7% 46.4% 45.8% 43.7% 15.0% 76.4% 41.0% 78.0% 63.68%
H2O0.9 74.60 28.13 21.69 25.41 52.83 24.55 41.73 33.95 15.23 22.51 27.64 69.77 19.13 -0.39%
SLM0.9 72.78 28.79 20.75 25.15 52.83 23.75 43.63 32.68 15.86 22.48 27.21 69.97 19.66 -0.59%
SnapKV0.9 70.20 27.90 20.08 25.38 52.72 24.04 43.93 33.92 15.56 22.49 27.65 70.58 19.05 -1.12%
PyramidKV0.9 75.44 28.79 24.46 24.72 52.16 23.64 43.39 32.10 14.11 22.22 23.70 70.48 19.05 -1.59%
CAKE0.9 74.23 27.23 23.16 22.87 51.18 21.32 43.78 34.65 14.20 21.19 22.33 69.22 19.22 -4.17%
H2O0.2 2.43 2.46 6.56 20.67 50.91 23.54 42.00 32.50 15.67 22.16 23.68 69.88 19.13 -23.33%
SLM0.2 3.21 3.35 8.82 18.02 40.11 20.12 36.45 29.91 15.01 21.11 24.68 66.17 18.69 -28.21%
SnapKV0.2 1.29 1.12 5.89 20.36 52.00 23.10 42.78 32.98 14.01 22.45 24.14 69.63 19.34 -22.89%
PyramidKV0.2 1.36 1.79 4.82 20.24 51.71 24.31 41.90 33.81 12.56 21.11 27.01 68.43 18.19 -25.33%
CAKE0.2 1.67 4.46 7.93 21.17 50.66 22.01 42.18 33.02 14.93 20.15 24.48 69.92 19.10 -23.47%
Mistral-7B-Instruct Full 33.36 29.24 6.83 20.82 39.68 28.74 37.80 33.87 22.88 22.19 22.94 86.87 16.08 100%
Oursk=1 31.54 29.02 6.71 20.81 39.80 27.20 38.93 33.46 22.15 22.39 22.67 86.81 15.33 -1.50%
Budget 89.6% 90.5% 84.2% 92.3% 84.1% 78.0% 97.9% 86.7% 74.4% 84.3% 89.1% 87.2% 89.4% 86.75%
H2O0.9 19.18 24.11 6.96 20.61 39.74 27.36 38.38 35.23 23.24 22.13 23.41 86.46 16.01 -4.29%
SLM0.9 31.16 27.68 5.89 19.88 39.15 27.26 37.66 35.06 21.94 22.31 21.73 86.63 13.67 -4.44%
SnapKV0.9 27.60 25.67 7.10 20.23 39.76 27.43 38.27 35.66 22.99 22.16 23.31 86.46 16.07 -1.90%
PyramidKV0.9 22.74 26.75 5.35 20.34 38.43 27.60 38.27 37.62 25.28 22.18 24.64 85.53 16.37 -3.15%
CAKE0.9 25.02 24.10 6.12 20.21 38.82 27.49 39.19 32.96 23.55 22.22 24.51 86.36 15.80 -4.14%
H2O0.2 1.52 4.46 4.02 18.45 35.55 27.28 33.45 34.36 23.21 21.80 20.93 85.99 14.82 -23.55%
SLM0.2 1.21 1.79 0.54 18.35 25.45 24.89 28.39 29.53 16.74 20.60 16.75 80.12 14.27 -35.58%
SnapKV0.2 1.44 0.20 3.35 17.81 37.44 26.76 36.02 33.90 23.05 22.13 20.88 84.27 16.20 -24.28%
PyramidKV0.2 1.14 0.58 1.40 19.24 31.20 26.19 34.53 37.25 24.55 22.24 22.18 86.58 15.76 -23.75%
CAKE0.2 4.55 1.03 2.09 18.21 30.07 26.25 36.03 32.39 22.79 21.53 23.96 86.61 15.00 -24.05%
Qwen2.5-7B-Instruct Full 88.02 31.70 29.85 24.55 61.43 20.81 43.17 47.15 30.70 23.64 24.24 87.64 2.44 100%
Oursk=1 88.02 31.25 30.39 24.48 60.01 20.66 42.74 47.20 29.56 22.64 24.04 87.65 3.58 +2.63%
Budget 90.7% 88.8% 86.5% 69.2% 84.1% 65.1% 69.2% 73.3% 64.4% 70.6% 85.4% 56.6% 84.4% 76.02%
H2O0.9 83.47 26.56 30.25 24.40 61.29 20.85 43.26 47.90 30.73 23.42 24.35 87.57 2.45 -1.46%
SLM0.9 88.93 30.80 28.25 24.30 60.91 19.86 42.54 46.02 28.75 23.06 23.42 87.03 3.05 -0.41%
SnapKV0.9 80.14 30.13 28.11 24.30 61.26 20.90 43.31 47.81 30.69 23.90 24.29 87.57 2.55 -1.01%
PyramidKV0.9 84.15 27.68 26.78 23.51 50.34 27.60 42.27 37.62 25.28 22.18 21.65 87.53 3.37 -2.74%
CAKE0.9 85.67 30.23 29.28 23.52 58.27 29.27 33.73 46.60 30.32 23.91 23.89 90.14 2.41 -0.06%
H2O0.2 5.91 3.57 17.40 21.70 55.39 21.41 40.30 45.84 29.72 23.04 20.84 87.99 2.15 -13.89%
SLM0.2 1.06 4.24 4.82 23.85 40.92 18.17 29.22 39.27 21.71 20.32 18.94 77.80 1.46 -31.66%
SnapKV0.2 2.58 7.37 15.13 21.52 57.06 21.40 40.93 46.96 29.87 23.58 20.69 87.10 2.44 -13.26%
PyramidKV0.2 9.70 8.57 9.60 26.42 51.19 26.42 39.48 39.54 25.73 21.02 21.65 85.31 1.76 -23.47%
CAKE0.2 7.58 9.11 5.56 23.66 49.19 30.20 43.04 46.35 29.00 23.33 22.94 85.45 2.89 -16.98%
Table 2: Performance of ReFreeKV and its comparison with five KV pruning methods on 13 datasets. Bold numbers indicate the best results aside from full-cache. Italics represent the real budget utilized by ReFreeKV. The correspondence between abbreviations and their full names of datasets can be found in Appendix. Avg. calculates the mean ratio of the model’s performance using different KV cache compression methods to its performance with the full cache. All average results (except for budget) are adjusted by subtracting 1 to provide a more intuitive understanding of the effectiveness of different methods.
3.1 Experimental Settings
Backbones
Our experiments are conducted with three LLM families of different model sizes: Llama3-Instruct with size of 8B/70B, Mistral-7B-Instruct-V0.3, and Qwen2.5-Instruct with size of 7B/32B/72B. We implement our ReFreeKV upon the released codebase of SnapKV***https://github.com/FasterDecoding/SnapKV. For the reduced attention matrix $A^{\prime}$, we set $k=1$ in practice (ablation provided in Section 3.4), and set $m=4$ for the initial cache ranking stage.
Datasets
For comprehensive evaluation, we evaluate ReFreeKV on datasets of both short and long context length of different domains, including mathematics, science, and commonsense reasoning on GSM8K DBLP:journals/corr/abs-2110-14168, GPQA DBLP:journals/corr/abs-2311-12022, TheoremQA DBLP:conf/emnlp/ChenYKLWMXWX23, TruthfulQA DBLP:conf/acl/LinHE22, and CoQA DBLP:journals/tacl/ReddyCM19. We also include tasks from Longbench DBLP:conf/acl/BaiLZL0HDLZHDTL24 with 8 datasets spanning document comprehension, summarization and coding. Appendix D provides a detailed description and statistics for all 13 datasets, along with how they are utilized in our experiments.
Evaluation Protocol
Our evaluation primarily assesses ReFreeKV ’s ability to preserve full-cache performance with its automatic pruning budgets. Accordingly, we compare ReFreeKV against its full-cache scores and also report the average compression ratio for each dataset.
Additionally, we compare with five prior KV cache pruning methods with varied budget sizes, including: Heavy Hitter Oracle (H2O) DBLP:conf/nips/Zhang00CZC0TRBW23, StreamingLLM (SLM) DBLP:conf/iclr/XiaoTCHL24, SnapKV DBLP:conf/nips/LiHYVLYCLC24, PyramidKV DBLP:journals/corr/abs-2406-02069 and CAKE DBLP:conf/iclr/QinCLHFCLL25. By evaluating performance consistency under fixed KV cache budgets of 90%, 50%, and 20%, we further illustrate the limitations arising from the input-specific threshold dependence.
Lastly, we also include a concurrent work Twilight DBLP:journals/corr/abs-2502-02770, which does not rely on a budget threshold but employs a top-$p$-inspired metric for adaptive token selection. It is worth noting that the $p$ value remains a hyperparameter to be tuned across models and inputs.
3.2 Main Results
The main experimental results are shown in Table 2. Due to space limitations, the results for the 50% budget size are placed in Appendix C. For comparison with Twilight, we separately report the results in Table 4 due to its different hyperparameter type. Based on Table 2, we can draw the following observations.
• ReFreeKV is able to fulfill our objective, capable of performing near-lossless dynamic compression across different models, varying input lengths, and diverse task types. Interestingly, with Llama3-8B and Qwen2.5-7B, ReFreeKV even surpasses the full-cache performance by 0.12% and 2.63% respectively, utilizing an average of 63.68% and 76.02% KV cache. With Mistral, ReFreeKV also manages to achieve near 15% compression with a relatively small 1.5% performance reduction. These results indicate that our proposed method can be suited for real-world scenarios with no bother by input-specific budget thresholds.
• In stark contrast, previous methods achieve a consistent full-cache performance only when manually determined a high budget ratio, e.g. 90%. However, when the budget is reduced, e.g. 50% and 20%, the degradation can become severe on certain datasets, distinct from ReFreeKV that automatically adjusts the pruning to always prioritize full-cache performance. Theoretically, one could tune the budget for each dataset that achieves minimal degradation, but this is generally infeasible for real-world open-domain instructions.
• ReFreeKV naturally reflects the difficulty of the generation task. As in Table 2, the dynamic budget ratio is high on Math&Science datasets (over 90%), while much lower on QA or Summarization datasets (as low as 15%). This observation is in line with our intuition, where inference on concise but hard tasks, such as math problems, requires more context and more precise calculation, resulting in higher budget allocation. From this aspect, our method design serves beyond for memory efficiency, but could be potentially leveraged for input analysis in a broader scope.
• Besides the dynamic compression, ReFreeKV also outperforms the three 90%-budget baselines, while itself uses less than 90% budget. Though, It is worth noting that the goal of this work is not to propose yet another KV cache pruning method that targets the best possible compression ratio under specific conditions. Instead, we seek to lift the threshold constraints and advocate for robust KV pruning that generalizes to arbitrary inputs. Through our proposed objective, we hope to further advance research in this direction.
GSM8K CoQA NarrativeQA Musique QMSum TriviaQA Avg.
Overall Prune Overall Prune Overall Prune Overall Prune Overall Prune Overall Prune Overall Prune
Llama3-8B-Instruct
Full 4.693 0.289 3.441 3.659 5.458 2.717 3.376
Oursk=1 4.638 0.034 0.331 0.033 3.159 0.255 3.658 0.264 5.330 0.241 2.684 0.211 3.300 0.173
H2O0.5 4.686 0.020 0.290 0.018 3.243 0.264 3.679 0.266 5.398 0.249 2.612 0.237 3.318 0.176
SLM0.5 6.071 0.006 0.301 0.030 3.230 0.257 3.784 0.259 5.454 0.234 2.628 0.208 3.578 0.166
SnapKV0.5 5.874 0.021 0.285 0.019 3.307 0.265 3.721 0.266 5.519 0.264 2.650 0.265 3.559 0.183
Llama3-70B-Instruct
Full 17.504 1.167 6.285 6.565 13.993 5.302 8.469
Oursk=1 15.975 0.154 1.253 0.290 6.042 2.345 7.412 2.352 13.835 3.028 5.224 1.602 8.290 1.623
H2O0.5 19.062 0.114 1.059 0.240 6.788 2.298 7.082 2.307 14.360 3.712 5.566 1.308 8.986 1.663
SLM0.5 22.079 0.150 1.138 0.229 6.804 2.069 7.107 2.053 14.734 3.700 5.576 1.367 9.573 1.595
SnapKV0.5 16.517 0.117 1.116 0.241 6.795 2.474 7.085 2.468 13.846 2.793 5.562 1.083 8.487 1.529
Table 3: The average inference time and pruning time of Llama3 with size of 8B/70B across six datasets, measured in seconds, with lower values indicating better performance.
Methods GSM8K NQA Qasper 2WQA Musique
Llama3-8B 75.28 24.06 43.91 35.33 14.77
+ Oursk=1 76.50 23.44 37.38 36.21 15.96
+ Twip=0.95 76.04 23.37 43.08 36.18 14.92
Mistral-7B 33.36 28.74 37.80 33.87 22.88
+ Oursk=1 31.54 27.20 38.93 33.46 22.15
+ Twip=0.85 30.33 27.17 38.90 33.23 22.92
Table 4: Performance comparison between ReFreeKV and Twilight across five datasets using Llama3-8B-Instruct and Mistral-7B-V0.3. Bold numbers indicate the best results aside from full-cache.
As separately reported in Table 4, we compare ReFreeKV with Twilight’s reported performance across the GSM8K, NarrativeQA, 2WikiMQA, and Musique datasets. Regarding the hyperparameter for Twilight, we adopt their reported optimal $p$ value: $p=0.95$ for Llama3-8B and $p=0.85$ for Mistral. As shown in the results, ReFreeKV achieves performance comparable to Twilight on both models. Notably, both methods frequently surpass the full-cache baseline across these datasets. Though, as Twilight requires different optimal $p$ values for each model, tuning may still be necessary. Nonetheless, both ReFreeKV and Twilight share the same principle of adaptive pruning, which we believe can further advance research in this area.
3.3 Efficiency Analysis
In this section, we conduct a quantitative study on the latency of ReFreeKV and the overall impact on the inference time, apart from the space reduction from KV cache pruning. We compare the time usage on six datasets with Llama3-8B/70B using ReFreeKV, along with three baselines with the 50% budget setting. Table 3 reports the time latency needed to apply pruning (Prune), as well as the average generation time after the prefilling stage of each sample (Overall).
From the results, the latency by ReFreeKV is on par with prior pruning methods. Notably, ReFreeKV achieves the best performance in 8 out of 12 comparisons of overall generation time, suggesting that the total generation speed of ReFreeKV is advantageous. The trend also remains consistent across different model sizes, underscoring the efficiency aspect of ReFreeKV.
For batch process, we further conduct analysis on throughput with different batch sizes. The full details of which are provided in Appendix F. The results confirm that ReFreeKV improves throughput upon the naive generation by 10-20%, especially it retains performance edges and robustness with increasing batch sizes.
3.4 Ablation Studies
We conduct ablation studies to investigate the impact of various configurations of ReFreeKV. We use Llama-3-8B-Instruct and perform experiments across five datasets, with results presented in Table 5. Appendix E provides additional results and analysis from our ablation studies.
Figure 3: Performance vs. Efficiency Trade-off. (a) Performance retention across five datasets (solid lines). (b) Computational budget consumption (dashed lines) relative to the dense baseline. The shared legend indicates the datasets. Results show that setting the universal threshold to $1\%$ could well balance between performance and memory budget, as it maintains robust full-cache performance while substantially reducing KV cache.
Methods GSM8K CoQA NQA Musique QMSum
Full 75.28 52.74 24.06 14.77 22.27
Post=1% Performance Using Different $k$
$k=1$ 76.50 52.86 23.44 15.96 21.95
Budget 93.2% 81.5% 48.7% 43.7% 15.0%
$k=1\%n$ 76.19 52.87 23.17 15.40 21.94
Budget 95.1% 94.8% 77.3% 77.2% 59.5%
$k=5\%n$ 75.59 52.75 21.62 13.71 21.43
Budget 98.6% 96.2% 45.1% 30.0% 33.1%
$k=10\%n$ 76.72 52.85 9.74 13.67 21.11
Budget 96.3% 96.8% 32.0% 19.7% 29.3%
$k=1$ Performance with Different Ranking Method
Attnt=1% 2.35 43.68 13.37 9.58 17.62
Budget 20.0% 9.7% 6.4% 6.4% 6.4%
Attnt=0.01% 54.66 51.58 17.45 14.79 20.90
Budget 61.7% 28.6% 6.7% 7.9% 7.1%
Table 5: Performance comparison for ablation studies in Section 3.4, with different $k$ values for attention matrix reduction and initial ranking alternatives.
Attention Matrix Reduction
The reduced attention matrix $A^{\prime}$ in Section 2.2 aggregates attention scores from the last $k$ rows. The upper part of Table 5 illustrates the model’s performance and the actual budget when setting $k=1$, $1\%n$, $5\%n$, and $10\%n$ ($n$ being the number of input tokens). It is clear that setting $k$ as $1$ achieves a significantly reduced budget, thus a higher compression ratio, with almost no change in performance compared to $1\%n$ and $5\%n$. On the other hand, while $10\%n$ can compress more KV cache, it fails to maintain performance (for instance, on NarrativeQA, the former achieves a performance of 9.74 using 32% of the budget, whereas the latter scores 21.44 using 48.7% of the budget). What’s even better is that since $k=1$ requires the least amount of computation, relying solely on the scores from the last row, the complexity of obtaining $A^{\prime}$ becomes $O(1)$, independent of the sequence length. The advantages of both high efficacy and low overhead make $k=1$ a solid design choice for calculating the norm metric.
Initial Ranking Strategies
Apart from the position-based ranking described in Section 2.2, we also investigate other alternatives, such as ranking by each token’s average attention score received from other tokens, similar to previous approaches such as H2O. The results are shown in the bottom of Table 5: under the same experimental settings, the attention-based ranking struggles to maintain a robust final performance. Empirical validation supports that position ranking is an appealing choice, offering both superior efficacy and efficiency.
The Universal Threshold
As we adopt the universal threshold as 1% for the attention norm difference, we comprehensively study the effects of smaller or larger thresholds, as shown in Figure 3. Intuitively, a larger threshold allows for a higher compression ratio but potentially more degradation, while a smaller threshold does the opposite. It is clear that when the threshold is set to a smaller value, such as 0.1%, the average budget increases as expected, yet the model’s performance sees little improvement. Conversely, when the threshold is set to 10%, it prunes more KV cache, but the model’s performance significantly deteriorates. Thus, we deem 1% as a reasonable and robust threshold for universal uses across models and tasks.
Generalization
We further conduct experiments to examine whether our design and hyperparameters can be generalized to LLMs of more sizes. On five datasets in Appendix Table 12, ReFreeKV with Llama3-70B and Qwen2.5-32B/72B demonstrates consistent near full-cache performance with the same ReFreeKV setting. Notably, the averaged compression ratio increases on datasets of longer context, achieving nearly a 50% compression.
4 Conclusion
In this study, we introduce a new KV cache compression objective that lifts the threshold dependency, so to achieve input-insensitive pruning for robust inference performance. Towards this objective, we propose a novel method, termed ReFreeKV, which employs a straightforward yet effective two-stage KV cache pruning process. Comprehensive experiments conducted across diverse datasets, encompassing a variety of tasks and context lengths, demonstrate that ReFreeKV achieves nearly lossless compression without input-specific tuning or thresholds, while with notable memory reduction.
Limitations
The main limitation of ReFreeKV stems from the gap between its current compression budget and the true optimal budget. This can be seen in Table 2 that for certain scenarios, e.g. QMSum with Mistral-7B, ReFreeKV reaches 84.3% budget while 50% budget is also viable with no performance degradation. We regard this gap as room for improvement in future work, facilitating more aggressive KV cache compression while ensuring the full-cache objective.
Another limitation of ReFreeKV is that though it demonstrates almost lossless compression, we rely on empirical experiments, and there is no hard guarantee on the degree of degradation. In Table 2, while both Llama3-8B and Qwen2.5-7B even surpass the full-cache performance, ReFreeKV with Mistral-7B shows trivial degradation of 1.5%. More robust methods could be developed as future work to further strengthen the robustness.
Appendix A Full Algorithm of ReFreeKV
Algorithm 1 ReFreeKV
Input: Prompt, Threshold $t$
Output: Compressed KV Cache
 Create Empty List $K_{c},V_{c}$
for Transformer Layer $L_{i}$ in LLM do
  $Q^{i},K^{i},V^{i}\leftarrow L_{i}(\text{Prompt})$
  $R^{i}\leftarrow$ Postion-Based Importance Rank
  $A_{last}^{i}\leftarrow\operatorname{Attention}(Q^{i}[\dots,-1,:],K^{iT})$
  $F_{b}^{i}\leftarrow\operatorname{Frobenius}(A_{last}^{i})$
  $A_{last}^{i}\leftarrow\operatorname{Square}(A_{last}^{i})$
  Reorder $A_{last}^{i}$ by Rank $R$
  $A_{cumsum}^{i}\leftarrow\operatorname{Cumsum}(A_{last}^{i})$
  $A_{cumsum}^{i}\leftarrow\operatorname{Sqrt}(A_{cumsum}^{i})$
  $A_{ratio}^{i}\leftarrow(F_{b}^{i}-A_{cumsum}^{i})/F_{b}^{i}$
  $\text{Index}^{i}\leftarrow\operatorname{Max}(\operatorname{Where}(A_{ratio}^{i}<=t))$
  $K_{c}^{i}\leftarrow\operatorname{Compress}K^{i}\text{by}R^{i}[I:]$
  $V_{c}^{i}\leftarrow\operatorname{Compress}V^{i}\text{by}R^{i}[I:]$
  Append $K_{c}^{i},V_{c}^{i}$ to $K_{c},V_{c}$
end for
return $K_{c},V_{c}$
Appendix B Retain the first two layers of LLM
In this section, we demonstrate the design choice of ReFreeKV to compress the KV cache starting from the third layer of LLMs. We apply ReFreeKV with Llama2-7B-Chat and conduct experiments on GSM8K and CoQA. We first perform a case study, followed by a comparison of the effects of not retaining full KV cache of the first two layers.
As shown in Table 6, when the threshold is set to 1% and no layers are frozen, the outputs for two examples on GSM8K and CoQA are incorrect and lack logical coherence. However, the budget exceeds 90% in both cases. By observing the actual budget of each layer, we can see that for these two examples, the budgets for layer 1 and 2 are relatively low, while the budgets from layer 3 to layer 31 are the same and very high, with the last layer being low again. We analyze that since the model, in the first two layers, is not yet able to identify truly important tokens, the attention distribution is relatively uniform. This may lead to the model discarding truly important tokens during eviction, resulting in subsequent generations being unable to access this information, ultimately causing the output to fail.
Based on the above case study, we attempt to retain certain layers of the model and explore the optimal configuration. We continue to test the results of retain different layers. The results are shown in Table 7. We can observe that freezing the first two layers achieves a balance between model performance and budget. Moreover, retaining the 31st layer does not significantly enhance the model’s performance and instead leads to an increase in budget. Consequently, ReFreeKV ultimately opts to begin KV cache compression from the 3rd layer of the model.
GSM8K
Input
A robe takes 2 bolts of blue fiber a-
nd half that much white fiber. How
many bolts in total does it take?
Budget
Layer 1: 67.71
Layer 2: 82.29
Layer $\sim3$1: 95.83
Layer 32: 37.50
Avg.: 93.90
Output
I have determined by answering the
answer to the format of bolts bolts b-
olts…(repeat)
Ground-Truth 3
CoQA
Input
You are given a story and a question.
Answer the question as concisely as
you can…Question: What color was
Cotton?
Budget
Layer 1: 42.14
Layer 2: 42.54
Layer $\sim3$1: 99.19
Layer 32: 79.64
Avg.: 95.03
Output
Question: Question: What is the que-
stion: What is the question:
Groud-Truth White
Table 6: Case Study.
Datasets Layer None 0 0,1 0,1,2 0,1,2,3 0,1,31
GSM8K Oursk=1 0.014 0.250 0.264 0.252 0.264 0.252
Budget 93.1% 97.0% 95.4% 97.5% 97.7% 98.2%
CoQA Oursk=1 1.47 52.80 53.58 53.58 53.32 53.46
Budget 92.7% 94.7% 95.5% 98.3% 98.4% 99.2%
Table 7: Performance of Llama2-7B-Chat on GSM8K and CoQA using ReFreeKV with different frozen layers.
Appendix C Main Results with 50% Budget
  Methods Math & Science CR Single-Doc QA Multi-Doc QA Summarization FSL Code   Avg.
GSM8K
GPQA
TheoQA
ThQA
CoQA
NrtvQA
Qasper
2WkMQA
Musique
QMSum
M-News
TriviaQA
Lcc
Llama3 Full 75.28 29.02 21.29 25.59 52.74 24.06 43.91 35.33 14.77 22.27 27.37 70.26 19.16 100%
Oursk=1 76.50 30.13 23.03 26.09 52.86 23.44 37.38 36.21 15.96 21.95 27.88 64.24 19.31 +0.12%
Budget 93.2% 86.7% 92.8% 78.7% 81.5% 48.7% 46.4% 45.8% 43.7% 15.0% 76.4% 41.0% 78.0% 63.68%
H2O0.5 31.08 2.90 16.47 17.99 42.37 23.14 43.05 32.79 15.77 22.79 26.24 69.83 19.18 -17.63%
SLM0.5 3.41 6.03 8.84 18.75 44.68 20.94 37.73 31.20 15.29 21.80 25.93 67.79 19.12 -24.73%
SnapKV0.5 12.96 7.81 16.87 19.52 42.84 23.40 43.93 33.98 15.94 22.67 26.07 69.66 19.19 -17.03%
Mistral Full 33.36 29.24 6.83 20.82 39.68 28.74 37.80 33.87 22.88 22.19 22.94 86.87 16.08 100%
Oursk=1 31.54 29.02 6.71 20.81 39.80 27.20 38.93 33.46 22.15 22.39 22.67 86.81 15.33 -1.50%
Budget 89.6% 90.5% 84.2% 92.3% 84.1% 78.0% 97.9% 86.7% 74.4% 84.3% 89.1% 87.2% 89.4% 86.75%
H2O0.5 2.50 8.26 5.89 20.28 38.53 27.62 36.97 34.31 23.17 22.32 22.60 86.52 16.45 -14.31%
SLM0.5 2.43 6.47 1.14 18.27 32.76 24.91 33.80 32.73 20.68 21.48 18.54 86.45 13.58 -27.61%
SnapKV0.5 3.03 8.93 6.83 19.03 39.22 27.15 38.27 34.60 22.94 21.93 22.68 86.31 16.18 -13.41%
Qwen2.5 Full 88.02 31.70 29.85 24.55 61.43 20.81 43.17 47.15 30.70 23.64 24.24 87.64 2.44 100%
Oursk=1 88.02 31.25 30.39 24.48 60.01 20.66 42.74 47.20 29.56 22.64 24.04 87.65 3.58 +2.63%
Budget 90.7% 88.8% 86.5% 69.2% 84.1% 65.1% 69.2% 73.3% 64.4% 70.6% 85.4% 56.6% 84.4% 76.02%
H2O0.5 34.12 14.73 20.88 23.10 59.69 21.59 43.37 47.60 30.81 21.00 22.95 85.54 2.24 -13.47%
SLM0.5 3.26 1.34 10.58 23.75 49.87 19.29 36.45 42.42 25.33 21.24 22.77 87.33 3.40 -23.56%
SnapKV0.5 20.77 12.28 22.76 22.25 60.21 20.93 43.22 47.45 30.61 23.76 22.70 87.57 2.21 -14.39%
Table 8: Performance of ReFreeKV and its comparison with five KV cache pruning models on 13 datasets. The cache size is set to 50%.
We show peformance comparison between ReFreeKV and different eviction-based methods with 50% cache size. The results are show in Table 8. From the table, we can observe the same conclusions with Section 3.
Methods GSM8K CoQA NQA Musique QMSum
Full 75.28 52.74 24.06 14.77 22.27
Post=1% Performance Using Different $k$
$k=1$ 76.50 52.86 23.44 15.96 21.95
Budget 93.2% 81.5% 48.7% 43.7% 15.0%
$k=1\%n$ 76.19 52.87 23.17 15.40 21.94
Budget 95.1% 94.8% 77.3% 77.2% 59.5%
$k=2\%n$ 76.19 52.82 22.87 14.34 21.74
Budget 96.0% 96.1% 69.8% 62.0% 47.1%
$k=3\%n$ 75.82 52.80 23.03 13.61 21.56
Budget 98.6% 95.7% 60.3% 46.0% 39.5%
$k=4\%n$ 75.21 52.76 22.96 13.45 21.47
Budget 98.8% 95.8% 51.9% 35.2% 35.4%
$k=5\%n$ 75.59 52.75 21.62 13.71 21.43
Budget 98.6% 96.2% 45.1% 30.0% 33.1%
$k=6\%n$ 75.66 52.81 21.95 14.33 20.94
Budget 98.5% 96.6% 40.4% 25.8% 31.6%
$k=7\%n$ 76.57 52.88 21.18 13.66 20.94
Budget 98.0% 96.7% 37.3% 22.5% 30.6%
$k=8\%n$ 76.35 52.86 20.36 13.75 21.11
Budget 97.4% 96.7% 35.1% 20.9% 30.1%
$k=9\%n$ 76.65 52.84 19.99 13.66 20.76
Budget 96.8% 96.8% 33.4% 19.9% 29.5%
$k=10\%n$ 76.72 52.85 9.74 13.67 21.11
Budget 96.3% 96.8% 32.0% 19.7% 29.3%
Table 9: Performance comparison with different $k$, expanded from Table 5.
BS=1 BS=2 BS=8 BS=16
[Input, Output] [4K, 8K] [8K, 16K] [512, 4K] [4K, 8K] [512,512] [4K, 4K] [512, 512] [2K, 2K]
Latency (s/100tokens)
HF Accelerate 3.61 5.29 3.50 5.26 3.48 12.23 4.71 13.01
Ours 3.19 4.08 2.84 3.60 3.06 10.42 4.25 10.64
Throughput (token/s)
HF Accelerate 27.71 18.89 57.08 38.01 230.00 65.41 339.43 122.98
ReFreeKV 31.35 24.51 70.30 55.56 261.83 76.81 376.89 150.40
Budget 73.3% 55.9% 62.0% 34.0% 76.7% 65.3% 82.8% 78.2%
Table 10: Performance comparison across various batch sizes and sequence lengths. We report latency (s/100 tokens, lower is better) and throughput (tokens/s, higher is better). In the header, ’k’ denotes 1024 tokens (e.g., 4K = 4096).
Appendix D Datasets Used in Experiments
In this section, we provide a comprehensive overview of all the tasks and datasets utilized in the experiments in this paper.
Math & Science
This task evaluates the model’s ability to tackle mathematical and scientific problems. By directly inputting questions and comparing the model’s output with the correct answers, we calculate the model’s Accuracy on these datasets: GSM8K is a dataset for evaluating model’s math-solving skills, featuring 8,000 elementary-level math word problems requiring basic arithmetic and reasoning. GPQA tests model’s understanding of physics concepts and problem-solving across various topics, assessing scientific reasoning abilities. TheoremQA evaluates model’s grasp and application of mathematical theorems, ranging from simple applications to complex proofs, testing advanced math skills.
Commonsense Reasoning (CR)
This task evaluates model’s ability to make deductions and understand everyday situations using implicit knowledge and logical inference. TruthfulQA (ThQA) evaluates model’s ability to generate accurate and truthful responses, testing models on distinguishing fact from fiction, especially in areas prone to misconceptions. We use BLEU as the metric. CoQA assesses model’s ability to understand and respond to questions in a conversational context, focusing on maintaining coherence and context throughout a dialogue. We use F1 Score as the metric.
Single Document QA (Single-Doc QA)
This task assesses the model’s reading comprehension skills when dealing with a single, extended document. NarrativeQA DBLP:journals/tacl/KociskySBDHMG18 is a dataset designed to evaluate model’s ability to comprehend and answer questions based on narrative texts, focusing on understanding stories and their underlying themes. Qasper DBLP:conf/naacl/DasigiLBCSG21 is a dataset aimed at assessing model’s capability to extract and answer questions from academic papers, emphasizing understanding complex scientific information. We employ F1 Score as the metric for above two datasets.
Multi-Document QA (Multi-Doc QA)
This task evaluates the model’s reading comprehension capabilities across multiple extended documents. 2WikiMultiHopQA (2WKMQA) DBLP:conf/coling/HoNSA20 is a dataset designed to test model’s ability to perform multi-hop reasoning and answer complex questions using information from multiple Wikipedia articles. MuSiQue DBLP:journals/tacl/TrivediBKS22 evaluates model’s skill in integrating and reasoning over information from multiple sources to answer comprehensive questions accurately. We leverage F1 Score as the metric for above two datasets.
Summarization
This task examines the model’s ability to comprehend and summarize lengthy documents. QMSum DBLP:conf/naacl/ZhongYYZMJACLQR21 is a dataset for evaluating model’s ability to generate concise summaries of meeting transcripts, focusing on capturing the key points from multi-party discussions. Multi-News (M-News) DBLP:conf/acl/FabbriLSLR19 is a dataset that challenges models to create coherent summaries by synthesizing information from multiple news articles on the same topic. We use Rouge-L as the metric for above two datasets.
Few-Shot Learning (FSL)
This task assesses the model’s few-shot learning capabilities. TriviaQA DBLP:conf/acl/JoshiCWZ17 is a dataset designed to assess model’s ability to retrieve and answer questions based on large collections of trivia, emphasizing comprehension and factual recall. We use F1 Score as the metric.
Code
This task evaluates the model’s ability to complete and generate code. LCC DBLP:conf/icml/GuoXD0M23 is a dataset focused on evaluating models’ ability to understand and generate code by considering extended code contexts, enhancing the ability to reason over complex programming structures. We use Edit Sim as the metric.
Appendix E Ablation
In this section, we present additional ablation study results for Section 3.4. By setting various values for $k$, we expand upon the results shown in Table 5. The expanded results are displayed in Table 9. These experiments facilitate a deeper understanding of how different parameter settings impact model performance and provide a basis for optimizing parameter selection.
As shown in Table 9, setting $k=1$ not only conserves pruning time but also achieves better model performance with a reduced budget.
Appendix F Analysis of Latency and Throughput
We perform following experiments for a comprehensive analysis of the latency and throughput. We follow the setup in FastGen using inputs of NarrativeQA, and conduct end-to-end latency comparison experiments on a single 80GB A100 GPU, using the standard HuggingFace (HF) Accelerate library as the baseline. The results are shown in Table 10.
GSM8K GPQA CoQA NrtvQA QMSum TriviaQA
Full 75.28 29.02 52.74 24.06 22.27 70.26
Oursk=1 76.50 30.13 52.86 23.44 21.95 64.24
Budget 93.2% 86.7% 81.5% 48.7% 15.0% 41.0%
SLM0.9 72.78 28.79 52.83 23.75 22.48 69.97
SLMF0.9 76.42 29.24 52.89 22.95 22.47 69.10
SLM0.5 3.41 6.03 44.68 20.94 21.80 67.79
SLMF0.5 4.55 3.13 44.59 21.31 21.88 68.04
SLM0.2 3.21 3.35 40.11 20.12 21.11 66.17
SLMF0.2 1.21 1.12 38.44 19.52 20.61 65.69
Table 11: Performance comparison between ReFreeKV and StreamingLLM with the first two layers frozen (SLMF) under difference KV cache budgets ($0.9$, $0.5$ and $0.2$).
Methods GSM8K CoQA NQA Musique QMSum
Llama3-70B-Instruct
Full 89.69 60.36 27.15 29.31 22.52
Oursk=1 89.76 60.38 26.94 28.88 22.30
Budget 93.3% 77.2% 58.7% 60.8% 52.9%
Qwen2.5-32B-Instruct
Full 91.36 58.03 24.78 40.04 22.84
Ours 91.81 57.29 22.65 40.54 22.52
Budget 91.6% 85.0% 68.2% 74.4% 76.8%
Qwen2.5-72B-Instruct
Full 90.22 54.14 24.36 42.13 23.93
Ours 90.30 54.19 24.10 41.70 23.31
Budget 90.9% 87.6% 63.4% 65.0% 68.6%
Table 12: Performance of LLMs with scales of 70B, 34B, and 72B across five datasets using ReFreeKV.
From the table, we can observe that ReFreeKV consistently outperforms the baseline in both latency and throughput. Moreover, ReFreeKV holds its advantage when the batch size increases, and its throughput remains robust. It is noteworthy that Table 4 in Section 4.2 also demonstrates the trivial overhead of ReFreeKV due to its simple matrix computation during its two-stage pruning process.
Appendix G Comparison with StreamingLLM
We perform additional experiments on six datasets using StreamingLLM (with Llama3-8B) that has the first two layers frozen, as an ablation study on the effects of the proposed pruning and halting process. As shown in Table 11, despite having the first two layers frozen, StreamingLLM still falls short of achieving the objective proposed in this work, not able to remain optimal consistently across datasets like GSM8K and GPQA as the budget varies. It is demonstrated that solely leveraging the position bias as in StreamingLLM is not able to maintain the full performance; the proposed pruning and halting mechanism that integrates the attention norm metric is necessary in achieving the goal for lossless pruning.
Appendix H Related Work
KV Cache Compression
To mitigate the large memory footprint of the KV cache during LLM inference, a prominent line of work has focused on pruning, which selectively discards less important tokens. Methods like Scissorhands DBLP:conf/nips/LiuDLWXXKS23 and SnapKV DBLP:conf/nips/LiHYVLYCLC24 retain tokens based on high attention scores. Others employ strategies based on token recency and historical importance, such as H2O DBLP:conf/nips/Zhang00CZC0TRBW23 or StreamingLLM DBLP:conf/iclr/XiaoTCHL24. FastGen DBLP:conf/iclr/Ge0LZ0024 further refines this by adapting retention strategies on a per-head basis.
A common limitation of these pruning methods is their reliance on a pre-defined memory budget, which can lead to inconsistent performance across different domains and datasets. In contrast, our work lifts this constraint, aiming for near full-cache performance with a dynamically sized budget without requiring task-specific tuning.
Beyond pruning, other compression paradigms have been explored. These include quantization-based methods, which reduce the numerical precision of cached values DBLP:journals/corr/abs-2405-06219; DBLP:conf/nips/HooperKMMSKG24; DBLP:journals/corr/abs-2503-18773; DBLP:journals/corr/abs-2402-12065; DBLP:journals/corr/abs-2502-04420, and low-rank approximation techniques that apply matrix decomposition to compress the KV cache DBLP:conf/iclr/ChangLLCHWHCAW25; DBLP:journals/corr/abs-2405-04434.
Attention Patterns
Our approach is grounded in research on the structural patterns of attention mechanisms. Studies have shown that attention heads often specialize, with some exhibiting predictable, position-biased behaviors such as consistently focusing on adjacent tokens DBLP:journals/corr/abs-1911-02150. This indicates a degree of redundancy in their learned patterns. Furthermore, the importance of individual heads can vary significantly across tasks DBLP:journals/corr/abs-2404-07965, suggesting opportunities for dynamic optimization.